### v1

# Rhetorical XAI: Explaining AI’s Benefits as well as its Use via Rhetorical Design

Abstract.
This paper explores potential benefits of incorporating Rhetorical Design into the design of Explainable Artificial Intelligence (XAI) systems. While XAI is traditionally framed around explaining individual predictions or overall system behavior, explanations also function as a form of argumentation, shaping how users evaluate system perceived usefulness, credibility, and foster appropriate trust. Rhetorical Design offers a useful framework to analyze the communicative role of explanations between AI systems and users, focusing on: (1) logical reasoning conveyed through different types of explanations, (2) credibility projected by the system and its developers, and (3) emotional resonance elicited in users. Together, these rhetorical appeals help us understand how explanations influence user perceptions and facilitate AI adoption. This paper synthesizes design strategies from prior XAI work that align with these three rhetorical appeals and highlights both opportunities and challenges of integrating rhetorical design into XAI design.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing HCI theory, concepts and models
1. Introduction
Modern AI systems are notoriously opaque, limiting efforts to understand or audit their behaviors (Cambria et al., 2023a; Saeed and Omlin, 2023). In response, Explainable Artificial Intelligence (XAI) aims to foster trust and accountability in AI systems by making their decision-making processes more transparent (Busuioc, 2021; Jacovi et al., 2021). XAI encompasses techniques to support human understanding of both local predictions (individual decisions) and global behaviors (Mohseni et al., 2021; Cambria et al., 2023b; Cabitza et al., 2023; Lai et al., 2023).
Different stakeholders leverage explanations to pursue distinct goals (Kim et al., 2024; Hoffman et al., 2021; Preece et al., 2018; Tomsett et al., 2018; Mohseni et al., 2021). For example, developers employ explanations to improve predictive performance by debugging model errors and biases (Hohman et al., 2018; Molnar, 2020). Ethicists and policymakers rely on explanations to assess the safety and accountability of AI systems (Busuioc, 2021). HCI researchers investigate how explanations can foster user appropriate trust (Zhang et al., 2020; Lai et al., 2023; Vasconcelos et al., 2023) and support complementary human-AI teaming (Bansal et al., 2020). Stakeholders with diverse objectives tend to evaluate explanations through different lenses, shaped by their specific roles and needs. In particular, end-users tend to judge explanations alongside more pragmatic and affective dimensions such as utility, trustworthiness, and alignment with personal values (Vera Liao et al., 2022). Therefore, a faithful and logically accurate explanation that neglects human-centered considerations may be disregarded altogether.
In this paper, we argue that designing explanations for effective adoption motivates expanding the scope of XAI beyond explaining how AI works towards articulating why AI merits use. An AI system is just one of many possible solutions to a user problem and AI explanations can serve a rhetorical function by communicating why such a system is beneficial, credible, and appropriate in a given context. Applications might include supporting critical evaluations of AI decisions (Wang and Yin, 2022; Bansal et al., 2020; Nguyen et al., 2018), assessing system credibility (Binns et al., 2018; Dodge et al., 2019), and calibrating appropriate user trust (Markus et al., 2021; Liao and Sundar, 2022).
Toward this end, we bridge Rhetorical Design with XAI. Rhetorical design provides an analytical framework to analyze both the how and the why of explanatory goals through three dimensions: (1) logical reasoning, how different explanation types appeal to distinct forms of reasoning; (2) credibility, how the AI system and its developers convey trustworthiness; and (3) emotional resonance, how different affective reactions elicited by explanations influence user perception and engagement with the system. Guided by this framework, we use a deductive-inductive approach to select, code, and analyze design examples from prior XAI research. Through this process, we synthesize design strategies across the three rhetorical dimensions. These identified strategies are situated in two explanatory goals: (1) How AI works, and (2) Why AI merits use.
Our paper offers both conceptual and practical contributions to the design of XAI systems. Conceptually, we propose a rhetorical framework for XAI that expands the traditional focus on explaining how AI works to include why AI merits use. Practically, guided by this framework, we analyze existing design strategies through the three classical rhetorical appeals (logos, ethos, and pathos), highlighting how explanations can be designed to be more logically convincing, credibly appealing, and emotionally engaging. By critically assessing the strengths and limitations of these strategies, we also discuss their broader implications for future XAI research and design.
This paper is organized as follows. In Section 2, we review key technical concepts and existing design paradigms in XAI, laying the foundation to highlight the missed opportunity of integrating rhetorical design into the current XAI landscape. In Section 3, we introduce and adapt rhetorical design into XAI through a conceptual framework. In Section 4, we describe how we apply our rhetorical framework to guide the qualitative coding process when reviewing prior XAI designs. We present our coding results in Section 5, elaborating on the identified design strategies. In Section 6, we critically examine the benefits and limitations of these strategies and offer recommendations for future XAI research. Finally, in Section 7, we summarize our contributions toward advancing rhetorical XAI.
2. Background
In this section, we briefly review key technical concepts (Section 2.1) and design paradigms (Section 2.2, 2.3, and 2.4) in XAI. This summary draws primarily from existing survey papers, position papers, and design frameworks in XAI, serving as background to contextualize our perspective on rhetorical design.
2.1. XAI Technical Concepts
Both AI and HCI researchers have acknowledged that making AI systems interpretable to people can foster more trustworthy and accountable use of AI (Markus et al., 2021; Molnar, 2020; Jacovi et al., 2021; Busuioc, 2021). Doran et al. (2017) categorize AI systems into three types — opaque, interpretable, and comprehensible — based on the level of transparency they offer in explaining automated decisions. In this taxonomy, interpretability refers to how well users can understand the internal processes by which an AI system transforms inputs into outputs. Other taxonomies of interpretability frame this input-output relationship through alternative lenses such as explanation scope, algorithmic techniques, or interactive format.
A common taxonomy based on the explanation scope is whether AI explanations are local or global (Molnar, 2020; Mohseni et al., 2021; Dwivedi et al., 2023). Local explanations explain why a specific input maps to an observed output. In contrast, global explanations provide an overarching view of how the AI generally maps inputs to outputs. Orthogonally, scope can also vary by model generality. Another foundational taxonomy differentiates model-specific techniques, which are tailored to a particular model, from model-agnostic techniques, which can be applied to any model.
Different XAI techniques lead to diverse types of explanations, which HCI researchers have evaluated with respect to their effects on users. These explanation types include example-based, feature-based, counterfactual, rule-based, and concept-based explanations (see Appendix A for further details). Each supports a distinct form of logical reasoning, reflecting the underlying algorithmic mechanisms that generate the explanation (Wang et al., 2019).
Given the wide range of human-AI interactions enabled by explanations (Bertrand et al., 2023; Xu et al., 2023; Chromik and Butz, 2021), many user studies adopt two primary forms of explanations — static vs. interactive — to examine how different explanation types influence user acceptance or foster appropriate trust in individual AI predictions. Static explanations typically present fixed outputs for given inputs and allow participants to engage with them over a limited number of trials. In contrast, interactive explanations enable users to actively engage with the system by adjusting inputs or model parameters, allowing them to explore alternative scenarios and gain a deeper understanding of the model’s behavior (Liu et al., 2021; Bertrand et al., 2023; Chromik and Butz, 2021; Nguyen et al., 2018).
While there is debate about the necessity of AI explanations, many researchers view XAI as a promising pathway for promoting safe, trustworthy, and accountable use of AI. For instance, when humans collaborate with AI to perform different tasks, explanations can help users understand model behavior, enabling them to judge when to accept or reject AI recommendations. This, in turn, fosters appropriate user trust and supports complementary human-AI teaming (Lee and See, 2004; Yang et al., 2020; Wang and Yin, 2021; Schemmer et al., 2023). To support cooperation, HCI researchers typically draw on human-centered principles when designing XAI systems (HCXAI). HCXAI emphasizes identifying user needs and contexts of use to inform whether, when, and how different explanations should be presented (Ehsan and Riedl, 2020; Ehsan et al., 2022b, c, 2024b).
Building on this perspective, the following sections briefly introduce primary XAI design paradigms including user-centered (Section 2.2), stakeholder-centered (Section 2.3), and sociotechnical approaches to explanation (Section 2.4).
2.2. User-centered Explanations
HCI scholars studying XAI have focused on how different types of explanations can support users in understanding AI systems more effectively (Hagras, 2018; Liao et al., 2020; Vera Liao and Varshney, 2021). Hagras (2018) first argued that XAI aimed to make AI understandable to humans. One strategy toward this goal is to reframe technical descriptions of AI decision-making into user-friendly language. For example, consider explaining how an autonomous car changes speeds. Instead of a technical explanation such as “if the distance to the car ahead is less than 2.5m and the road is 10% slippery then reduce car speed by 25%,” a more human-like natural language rephrase might be “if the distance to the car ahead is low and the road is slightly slippery, then slow down.” In general, researchers have developed various types of explanations (see appendix A) that employ visualizations and natural language to make AI more accessible and understandable to people (Lai et al., 2023; Robbemond et al., 2022).
Given the diversity of explanations, researchers have proposed different frameworks to help designers select the most appropriate explanation type for a given context. For example, Liao et al. (2020) proposed an XAI “question bank” comprised of common user inquiries such as How AI works, Why AI made a specific prediction, Why not an alternative prediction, What AI would predict if the input changes, and How to change the input to alter the AI prediction. Other researchers have adopted this question-informed framework to select AI explanations when designing XAI systems (Xin He and Zhang, 2023; Ehsan et al., 2022a; Wang et al., 2023a). Other frameworks aim to bridge human understanding and machine explanations from a cognitive perspective, such as Wang et al. (2019) proposing to design AI explanations align with human cognitive processes, Chen et al. (2022) incorporating task-specific human intuitions when assessing user understanding of explanations, and more (Zhang et al., 2019; Subhash et al., 2022; Chen et al., 2023a).
Since AI systems serve diverse end-user groups, explanations can be tailored to their specific needs and contexts (Ehsan et al., 2024a; Kim et al., 2024). For example, loan seekers and lenders have different explanatory needs for understanding how a credit scoring AI system functions (Bove et al., 2022). Loan seekers might be particularly interested in why their financial record led to a rejection and how to alter the decision. Lenders might seek a global understanding of AI behavior across historical loan records to support financial analysis, enabling them to make more informed decisions about issuing loans to future applicants (Dikmen and Burns, 2022). Similar distinctions are also prevalent in the medical domain between patients and healthcare providers (Burgess et al., 2023; Kim et al., 2024). Collectively, these cases highlight how domain expertise shapes explanatory needs (Mohseni et al., 2021; Kim et al., 2024). Lay users often prefer personalized and actionable guidance, while experts commonly seek system-level understanding.
2.3. Stakeholder-centered Explanations
While end-users are a central focus in XAI, they represent just one of the many stakeholders involved in developing, deploying, and adopting AI systems. To situate AI within this broader socio-technical ecosystem, explanations should also cater to stakeholders’ distinct goals, responsibilities, and information needs. Two primary approaches for categorizing stakeholders are based on their roles within the AI lifecycle (Tomsett et al., 2018; Preece et al., 2018; Hoffman et al., 2021; Dhanorkar et al., 2021) and their level of AI knowledge (Mohseni et al., 2021).
In role-based stakeholder categorization, Hoffman et al. (2021) demonstrated distinct explanatory needs between developers, policymakers, and domain users. They reported that developers mainly employed explanations to debug AI errors and identify potential algorithmic biases to improve overall predictive performance. Policymakers leveraged explanations to audit AI behavior in expected deployment settings and unexpected edge cases. Domain users used explanations to understand or counterfactually remediate unfavorable AI decisions. Other researchers also used role-based frameworks but with alternative groupings of stakeholders (Tomsett et al., 2018; Preece et al., 2018; Dhanorkar et al., 2021).
Complementary to stakeholder roles, Mohseni et al. (2021) proposed a taxonomy of explanation needs based on varying levels of AI expertise. They distinguished between AI novices (layman users) from AI experts (engineers and developers), each leveraging explanations to suit different purposes. However, Suresh et al. (2021) contended that both role-based and AI knowledge-based frameworks could be problematic, as overlapping explanation needs could challenge consistency between different stakeholder groups. Consequently, they proposed to organize stakeholder needs around Goals, Objectives, and Tasks along the AI development cycle, aiming to offer a more cohesive and ontologically grounded vocabulary.
While stakeholder frameworks might differ in structure and granularity, they all highlighted the need to contextualize AI explanations to specific goals and responsibilities of different stakeholder groups (Wolf, 2019; Carroll, 2003). Identified stakeholder needs then served as specifications to guide the technical development (Kaplan et al., 2024) and interface design of XAI systems (Dhanorkar et al., 2021; Ehsan et al., 2023).
2.4. Socio-technical Explanations
As researchers broaden their study focus from end-users to a wider range of stakeholders, this shift reflects a socio-technical perspective for studying XAI (Preece et al., 2018; Ehsan and Riedl, 2020; Dhanorkar et al., 2021). Brandão et al. (2019) argued that AI experts often lack a natural understanding to connect the potential social meanings with models they create. Ehsan et al. (2023) further illustrated this disconnection through a use case where an AI-powered sales prediction tool, despite its 90% reported accuracy, was adopted by only 10% of sellers due to either over-trust by less experienced users or disregard by experienced sellers. Similar cases have emerged in different areas, including customer-service relationship (Northey et al., 2022), journalist norms and values (Komatsu et al., 2020), and frontline workers (Fox et al., 2023).
The main issue described in the above use cases is that, current AI deployment did not account for organizational social needs (Ehsan et al., 2023). In particular, users reported that verifying AI predictions often created added burdens for organizational work, such as having to learn new terminologies, consult technical experts, or manage concerns about reduced transparency. To address this, some XAI researchers argued that explanations should also account for users’ roles, expertise, and collaborative practices within the organization (Ehsan et al., 2021, 2023).
2.5. Research Focus of this Study
Thus far, Section 2 has outlined key technical concepts in XAI, covering commonly used terminologies and taxonomies (2.1), as well as design paradigms of human-centered XAI, including user-centered (2.2), stakeholder-centered (2.3), and sociotechnical perspectives (2.4).
In this work, we expand the design paradigms of human-centered XAI to incorporate the rhetorical dimensions of user experience. While traditional XAI prioritizes functional utility and cognitive human-AI alignment across different stakeholder objectives, we argue that pragmatic and affective responses equally shape understanding and adoption of AI (Dewey, 2005; Buchanan, 1985a, 2001a). For example, when interacting with XAI systems, users frequently grapple with questions such as: Will this explanation help me achieve my goal better? Can I trust its rationale? Does it align with my ethical principles? These questions highlight that technical fidelity and cognitive alignment alone are insufficient; explanations must also resonate with users’ expectations, emotions, and values to be meaningful and persuasive.
Our reframing positions AI explanations not only as informative responses but also as rhetorical artifacts that mediate the relationship between AI systems and end-users. In other words, AI is just one of many possible solutions to a user problem. AI explanations thus must serve a persuasive and contextualizing role by communicating why an AI system is beneficial, credible, and appropriate in a given context. From this perspective, a rhetorical XAI design can more effectively support users in evaluating automated decisions (Wang and Yin, 2022; Bansal et al., 2020; Nguyen et al., 2018), assessing system credibility (Binns et al., 2018; Dodge et al., 2019), and calibrating appropriate user trust (Markus et al., 2021; Liao and Sundar, 2022).
In this paper, we investigate the argumentative role of AI explanations in mediating user adoption of AI systems and we synthesize how prior research has employed various design strategies to fulfill this rhetorical function. We draw on the concept of Rhetorical Design, a longstanding product and graphic design practice that examines how form, structure, and presentation influence user adoption. In the next section, we introduce the theoretical foundations and design practices of rhetorical design in HCI and describe how we adapt this lens to examine and inform explanation design.
3. Rhetorical XAI
Originally developed in product design (Buchanan, 1985a), rhetorical design has expanded into graphic and communication design (Sosa-Tzec and Siegel, 2014; Sosa-Tzec et al., 2015; Foss, 2004; Barnard, 2013; Boyarski and Buchanan, 1994), information visualization (Hullman and Diakopoulos, 2011; Hua, 2021), and other HCI domains (Wahyurini, 2021; Rutsky, 1999; Brummett, 1999). Specific rhetorical strategies include visual metaphors and analogy, emphasis and contrast, information alignment and proximity, and more.
In this section, we first introduce the original concept of rhetorical design in product design (Section 3.1) and outline common strategies (Section 3.2) employed across design fields. Building on this foundation, we establish rhetorical design as an analytical framework (Section 3.3) for identifying design strategies from prior XAI research.
3.1. Rhetorical Design
Daniel Kahneman, Nobel Prize winner in Economics, has shown through his research that people often make decisions in ways that deviate from strict rationality (Kahneman, 2012). In product design, Buchanan (1985a) argued that product quality is not solely determined by functional aspects that reflected technical capabilities, but also by the credible signals and emotional cues embodied in its design. This is particularly evident when people choose a product not just for its utility, but also for factors such as its identity (Clark, 2009) or broader social significance (Stoekl, 2009; Forlizzi, 2008; Liu and Kim, 2023).
What makes a product compelling? To answer this question, Buchanan (1985a) reviewed a list of everyday products — from household utensils to electronics — then drew an analogy between a product and a verbal argument. He argued that a product functions as a rhetorical artifact designed to compel interest, attract attention, and invite use from its intended audience. From this view, a product’s rhetorical appeal extends beyond meeting practical needs by inviting users to envision new ways of acting, thinking, and being. Here, Buchanan drew on classical rhetoric (i.e., logos, ethos, and pathos) to contend that product design, like persuasive speech, seeks to influence how users act, think, and feel.
In classical rhetoric, Aristotle (1926) argued that an argument is persuasive not only due to logical reasoning (logos) but also because of the speaker’s credibility (ethos) and emotional appeal to the listener (pathos). For example, a genuine diamond ring (strong logos) may appear counterfeit if offered by a financially impoverished individual (weak ethos). Likewise, simple truths (strong logos) cannot be readily accepted by a child crying and throwing a tantrum (weak pathos).
Extending these rhetorical principles to product design, Buchanan (1985a) introduced Design Rhetoric, where logos refers to the product’s technological reasoning, ethos to the implied character or credibility of the manufacturer as reflected in the product, and pathos to the product’s usability and emotional appeal for intended users; Here, technological reasoning is the intelligent organization of materials and processes to solve practical problems in ways that appear logical, efficient, and appropriate to users’ needs and values. Effective designs balance these three rhetorical appeals through a designer’s own judgment (Buchanan, 2007).
To summarize rhetorical design, two main points emerge: 1) in product adoption, design functions as rhetorical arguments with specific strategies to make the product persuasive and compelling for use; 2) these strategies can be framed around three unique dimensions – logical reasoning, credibility, and emotional resonance – each aligning with the key actors in the rhetorical triangle: the product, the designer, and the user.
3.2. Design Strategies for Different Rhetorical Appeals
HCI practitioners have developed various strategies to fulfill the three rhetorical appeals. Here, we briefly review examples from User Interface (UI) design and data visualization, as they closely resonate with design practices in the XAI community. In web design, Ibrahim et al. (2013) observed that visually salient features such as “contrast, liking, or surprise” are effective in evoking emotional responses from users, credibility is supported by elements like “recognizability, assurance,” or “evidence and social proof” by showing peer activities on the web, and logical appeal is reinforced through features such as “simplicity, variety, and verifiability.” Similarly, Sosa-Tzec and Siegel (2014) analyzed how visual strategies and affective heuristics operate in graphical interfaces to elicit these rhetorical effects. They later published a rhetorical design handbook tailored for UI/UX contexts (Sosa-Tzec et al., 2015).
In information visualization, Hullman and Diakopoulos (2011) outlined a variety of linguistic (e.g., irony, similarity) and visual (e.g., visual metaphors, analogy, contrast, classification) strategies that enhanced narrative storytelling. They presented Mapping America [^1] Mapping America: Every City, Every Block. The New York Times: http://projects.nytimes.com/census/2010/explorer and Poll Dancing [^2] Poll Dancing: How Accurate are Poll Predictions? Datablog: https://thesocietypages.org/graphicsociology/2010/06/03/poll-data-poll-dancing/ as examples to illustrate how different strategies cohesively interact to make presented information more logically convincing, credible, and emotionally engaging. Many designers employed these strategies in infographics and visualizations to communicate complex social issues. For example, Hua (2021) used data association, embodiment, and visual metaphor to design infographics to raise public awareness about recycling contamination. Similarly, Meuschke et al. (2022) applied narrative-driven strategies to facilitate public understanding of medical data. Other directives include supporting civic engagement and participatory discourse (Dörk et al., 2013; Claes and Vande Moere, 2017).
This paper aims to synthesize design strategies (similar to those discussed above) for crafting AI explanations. In the following section, we introduce the rhetorical framework for identifying and categorizing these strategies, which guided our qualitative coding and review process of design examples from prior XAI research.
Dimensions Rhetorical Inquires for AI Explanations
Explain How AI Works Explain Why AI Merits Use
Logical
Reasoning
What forms of explanations are used, and how are they presented to convey the underlying logic of AI local prediction or overall behavior? How do these explanations help end-users apply AI effectively to address their real-world needs and tasks?
Credibility How does the framing and presentation of explanations signal their credibility to users? How do the framing and presentation of explanations signal the credibility of AI systems?
Emotional Resonance How do the framing and presentation of explanations engage user emotions to facilitate their understanding of AI? How do the framing and presentation of explanations engage user emotions to better utilize the AI system?
Table 1. A Rhetorical Inquiry Framework to Characterize AI Explanations.
3.3. A Rhetorical Framework to Characterize Explanation Design
Our framework builds on two primary ideas of rhetorical design, as outlined at the end of Section 3.1. Drawing on the first idea — in product adoption, design functions as rhetorical arguments with specific strategies to make the product persuasive and compelling for use — we propose extending XAI beyond merely (1) explaining How an AI system works (e.g., clarifying local predictions and global behaviors) to also (2) explaining Why an AI system merits use.
Our call to broaden XAI resonates with prior work, which argued that effective use of AI may not require a deep understanding of model behaviors. For instance, Miller (2021) noted, “Explanations of how models work might just distract people from figuring out what they really need.” Regardless of the specific needs (see Section 2.2), from a user-centered perspective, understanding how AI works (i.e., pure interpretability) is means rather than an end for why people choose to use AI in the first place. Therefore, explanations should go beyond the current focus on functional transparency, addressing more rhetorical questions such as: What is the compelling value proposition for a user to engage with AI?
Next, we apply the second idea — the rhetorical triangle of logical reasoning, credibility, and emotional resonance — as qualitative coding axes to guide our synthesis of design strategies in prior XAI literature. This framing also resonates with recent calls emphasizing that user perceptions of AI are influenced by more than just logic (Thommes et al., 2024). Experiential elements, such as tangible features (Ghajargar and Bardzell, 2023; Alvarado et al., 2022), embodied interaction (El-Zanfaly et al., 2023), and sensory or emotional responses of the AI system (Bryan-Kinns, 2024), can play a more significant role in influencing users to accept or reject AI recommendations.
Inspired by this line of work, we ask what if designers viewed themselves as rhetoricians? That is, by drawing on the rhetorical triangle: (1) logical reasoning conveyed through different types of explanations; (2) credibility projected by the AI and its developers; and (3) emotional resonance elicited in users, how might we craft compelling design arguments about AI through explanations? What design strategies exist, or need to be developed, to fulfill each rhetorical appeal and strike a thoughtful balance among them?
Building on this conceptual adaptation, we argue that rhetorical design can serve two distinct objectives in XAI: (1) supporting traditional AI explanations that clarify How AI Works, and (2) designing explanations that communicate Why AI Merits Use to encourage adoption. We examine each objective through the three rhetorical dimensions: logical reasoning, credibility, and emotional resonance. In Table 1, we outline each rhetorical appeal as a coding inquiry across the two goals, which guided our open and axial coding process detailed in the next section.
4. Method
We conducted a narrative review guided by coding inquiries specified in Table 1. Unlike a systematic review, which follows established protocols such as PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) (Moher et al., 1996), we chose a narrative review to allow for greater flexibility in selecting, interpreting, and synthesizing diverse design strategies across different XAI contexts (Baumeister and Leary, 1997; Ferrari, 2015). Nevertheless, potential author bias remains an inherent limitation of this approach (see limitations in Section 6.3). To mitigate this, we employed a structured three-step process involving both open and axial coding.
Step 1 - Classifying Design Examples based on XAI Objectives: Before applying open coding to identify specific design strategies, we first conducted a top-down process to classify prior work based on whether AI explanations were primarily used to (1) convey how the AI system works or (2) enhance the usability and adoption of the AI. Specifically, we conducted a need-based meta-analysis grounded in the principles of user-centered XAI (Section 2).
For example, Cai et al. (2019a) presented example-based explanations designed to help users understand how an AI model recognizes a drawing of an avocado. Their primary motivation was to “help users develop intuition for the reasons underlying machine predictions, without directly exposing the internal logic of the algorithm.” Therefore, we classified Cai et al. (2019a) under the first XAI objective: explaining how AI works. In contrast, in another work by Cai et al. (2019b), the authors integrated example-based explanations and other interactive features into a case-based retrieval system to inform clinical diagnosis. We classified Cai et al. (2019b) under the second XAI objective since the presented explanations primarily demonstrate the practical benefits of the AI system. Additionally, we classified Wang et al. (2023b)’s interactive recourse tool for loan approval under both XAI goals, as explanations provided not only support user understanding of model behavior but also offer actionable insights to help users navigate the application and address real-world challenges.
We further validated our classification by examining the AI expertise of recruited participants and the relevance of the AI task to their professional work. As a rule of thumb, studies involving domain experts were more likely to pursue the second objective, since specialists could more easily evaluate the utility of XAI systems in real-world contexts. Conversely, crowdsourced studies were more likely to focus on the first objective, as the tasks were typically less grounded in the specific needs and contexts of real-world users. Of these, we flagged studies that employed attention checks or screening criteria to recruit participants with relevant domain knowledge.
Step 2 - Identifying Design Strategies: During open coding, we first collected and organized prior XAI systems based on different types of explanations (Appendix A). Next, we scrutinized the presentation of explanations, focusing on design components such as visual elements (e.g., lines, shapes, and graphics) and linguistic features (e.g., languages and tones). We inductively derived an initial set of codes characterizing recurring design strategies from this process.
Step 3 - Associating Design Strategies with Rhetorical Appeals: Finally, we conducted axial coding to organize recurring design patterns by aligning them with the three rhetorical appeals: logical reasoning, credibility, and emotional resonance. We analyzed reported quantitative and qualitative user feedback to assess how design strategies influenced user perceptions. This step connected design intent to user impact by examining how specific strategies contributed to rhetorical persuasiveness.
For example, for quantitative measurements, “simulatability” (i.e., participants predicting AI behavior) and “perceived usefulness” assess the effects of logical appeals, as they reflect improvements in user understanding of AI reasoning and AI utility. For credibility appeals, “self-confidence” and “perceived trust” indicate how trustworthy the AI system appears. Similarly, for emotional appeals, “affective reaction scores” and “user engagement metrics” quantify the extent to which explanations resonate affectively with users.
Additionally, we contextualized quantitative measures with reported qualitative feedback. We provide our reviewed articles and corresponding coding schemes in Appendix B. For each paper, we specified the primary XAI goal, identified design strategies, and mapped corresponding quantitative and qualitative feedback to appropriate rhetorical appeals.
However, we acknowledge that our meta-analysis is selective rather than comprehensive. For example, in some cases, design strategies identified from papers primarily targeting the first XAI goal could also be applicable to the second goal. Moreover, not all user studies in prior work captured responses across all three rhetorical dimensions. We further elaborate these limitations in Section 6.3.
5. Results
Drawing on our coding process (Section 4), we report synthesized design strategies from prior XAI research in this section. We categorize strategies based on three rhetorical appeals: Logical Reasoning (Section 5.1), Credibility (Section 5.2), and Emotional Resonance (Section 5.3). Within each appeal, we describe how different design strategies support the two XAI objectives: explaining how AI works or why AI merits use (see Table 1). We assess each strategy’s strengths and limitations based on our axial coding process, linking reported user experiences to their intended rhetorical appeals.
5.1. Logical Reasoning
A key focus of XAI is explaining the logical reasoning behind automated decisions to foster appropriate trust in AI predictions or align user mental models with overall AI behaviors (Wang et al., 2019; Miller, 2019; El-Assady and Moruzzi, 2022). Here, we identify three design strategies from prior XAI research: Isolated Presentation, Adaptive and Mechanistic Presentation, and Graphical Synergy. Explanations can also communicate why AI merits use by positioning the AI as an intelligent, efficient, and rational collaborator. Toward this end, we identify two design strategies that directly enhance the usability and adoption of an AI system: Explanations as Recommenders and Explanations as Retrievers.
5.1.1. Isolated Presentation
Different types of AI explanations inherently embody distinct forms of human logical reasoning. Wang et al. (2019) conceptually mapped how various machine learning models reason and generate explanations to corresponding forms of human decision-making. For example, they argued that Bayesian probability represents inductive reasoning, a bottom-up approach that uses prior intuitions to draw broader conclusions. Case-based models encapsulate analogical reasoning, which solves new problems by synthesizing solutions from previously solved similar issues. Rule-based models illustrate deductive reasoning, a top-down approach that begins with general principles and operationalizes them to specific cases.
Explanations can be compelling logical arguments when their embodied reasoning aligns with user mental models. When explaining how AI works, the presented explanation reinforces trust and interpretability if users perceive the mapping from inputs to outputs as logically coherent or aligned with their expectations. In summary, different types of AI explanations operationalize different modes of human reasoning, and their persuasiveness depends on how well embodied logic align with user expectations.
A dominant design strategy employed in prior XAI studies is isolated presentation, where different types of explanations are presented individually for A/B testing. Specifically, these studies examined the extent to which distinct logical structures embodied in different explanations (e.g., features, examples, rules, or concepts) helped to improve human-AI complementary performance (Bansal et al., 2020; Gajos, 2024; Nguyen et al., 2018) or promote appropriate reliance on AI predictions (Schemmer et al., 2023; Yang et al., 2020; Zhang et al., 2020; Liu et al., 2021; Vasconcelos et al., 2023). For example, in comparing feature-based and example-based explanations, Chen et al. (2023b) observed that the latter elicited more critical reflections on potential misalignments between human intuition and machine reasoning. Moreover, van der Waa et al. (2021) demonstrated that rule-based explanation showed a small positive effect on (global) system understanding compared to example-based explanation. Nevertheless, both types of explanations could lead to confirmation bias (Bertrand et al., 2022), where users tended to over-rely on AI predictions erroneously.
Regardless of format or style, all AI explanations are paraphrased distillations of the underlying computational algorithm. Isolated evaluations of explanations may simplify analysis, but can also mislead users by reinforcing prior (erroneous) assumptions or painting an incomplete picture of system behavior (El-Assady and Moruzzi, 2022). These risks arise in part because AI systems and the explainable techniques used to interpret them (described in Section 2.1) are not inherently designed to align with human reasoning (Chen et al., 2023a; Wang et al., 2019). Although neural networks could be loosely modeled on the structure of human brain (Schaeffer et al., 2022), AI systems are typically trained to rely on complex, uninterpretable features that differ fundamentally from the semantic reasoning humans use to understand the world (Fei-Fei et al., 2007; Zhang et al., 2019). As a result, explanations may appear plausibly persuasive to users, but do not necessarily reflect the actual decision-making process of the system or reliably align with human reasoning.
Furthermore, explanations may be misinterpreted when subtle variations in visual elements convey unintended or oversimplified meanings (Lupton and Phillips, 2015; Gibson, 1899). For example, Ye et al. (2023) observed that subtle differences in saliency maps design could significantly influence annotator perceptions in uncertainty-aware medical semantic segmentation. They compared three formats — singular highlights, continuous gradient maps, and confidence contour maps — and found that the first two were less interpretable. In particular, continuous gradients were prone to confusion, as annotators struggled to infer clear thresholds for distinguishing uncertainty levels. Other studies reported similar qualitative feedback, where participants noted that continuous gradient maps of local feature importance only helped them recognize relevant features but made it difficult to quantify their relative importance (Alqaraawi et al., 2020). These failures were also observed in textual contexts (Schuff et al., 2022).
5.1.2. Adaptive and Mechanistic Presentation
Since AI systems do not inherently follow human reasoning, presenting isolated explanations risks reinforcing misconceptions by oversimplifying AI behavior. To mitigate this, researchers have adopted two design strategies. Adaptive presentation dynamically adjusts explanation format based on evolving user objectives or deployment contexts. Mechanistic presentation combines multiple explanation types into an integrated UI dashboard. We discuss these two strategies in tandem because both extend isolated presentations through adaptive adjustments or mechanical aggregations. While the mode of delivery differs, the underlying logical appeals of these explanations remain unchanged. This contrasts with synergistic design strategies introduced later (Section 5.1.3), which fundamentally reconfigure explanatory logic.
Adaptive explanations are primarily delivered in two ways: (1) through user requests, such as in a conversational setting, or (2) automatically triggered through internal system indicators. For user-initiated requests, Miller (2019) showcased an idealized explanation dialogue in which an XAI agent provides different types of explanations (e.g., features, rules, examples) in response to diverse user inquiries on how the AI classifies an input image as a spider (versus a beetle or octopus). Shen et al. (2023) operationalized this adaptive dialogue in human-AI scientific writing contexts. They constructed a chatbot to provide heterogeneous AI explanations across different stages of human-AI communication.
For system-triggered deliveries, explanations are automatically adjusted in response to internal (invisible to end-users) indicators. For example, some XAI systems present explanations only when AI confidence is low (Bansal et al., 2020) or dynamically adjust explanations based on real-time accuracy updates or error boundaries (Bansal et al., 2019b, a; Wang and Yin, 2023). Others adapt explanations based on human learning progress (Wambsganss et al., 2020) or show explanations only when the AI’s predicted correctness is estimated to exceed that of the user (Ma et al., 2023).
Compared to adaptive presentations, mechanistic presentations of explanations are generally more straightforward to implement since no real-time adaptations are required. For example, Wang et al. (2019) integrated local feature importance, global feature contributions, and counterfactual rules into an AI-assisted medical diagnosis tool, enabling clinicians to consult multiple explanation types simultaneously when evaluating patient cases. Similarly, Bove et al. (2022) developed an explainable auto-insurance interface that helped vehicle owners understand how their driving behaviors influenced insurance premiums. Likewise, Cau and Spano (2025) created an explainable hiring dashboard that enabled recruiters to accept or reject applications based on applicant attributes presented through multiple concurrent types of explanations.
Overall, both adaptive and mechanistic presentations offer unique benefits to improve user understanding and support human-AI teaming. Adaptive presentations enhance relevance by tailoring explanations to user goals, expertise, or decision context, reducing cognitive overload and aligning responses with moment-to-moment needs. Mechanistic presentations offer breadth and transparency by juxtaposing multiple explanation types in a unified interface, allowing users to cross-reference local and global reasoning or compare factual and counterfactual views. Compared to isolated presentations, interactivity is a crucial differentiating factor, allowing logical reasoning processes to unfold dynamically in response to user actions or system feedback.
Despite their benefits, Gaole et al. (2025) found that both adaptive and mechanistic presentation formats could lead to a clear over-reliance on the AI system. The authors attributed this over-reliance to an illusion of explanatory depth where the presence of multiple explanations, regardless of actual fidelity, reduced critical evaluation of system behavior. This echoes a classic rhetorical strategy in communication, where shifting the burden of proof makes arguments appear more persuasive (Hahn and Oaksford, 2007).
5.1.3. Graphical Synergy
Regardless of whether explanations are presented in isolated, adaptive, or mechanistic formats, different explanation types often coexist unilaterally without being cohesively integrated. As Wang et al. (2019) noted, this fragmentation arises because the underlying explanatory algorithms rely on fundamentally different logical bases. Nevertheless, emerging work has shown promise in organically merging diverse explanation types into a visually coherent representation, an approach we term graphical synergy.
For example, Evans et al. (2022) synergistically combined normative example-based explanations with counterfactual explanations, forming a cohesively contrastive causal chain. Specifically, they presented prototypical examples alongside counterfactual instances of cell nuclei to help clinicians differentiate between positive and negative Ki67 scores used to estimate tumor proliferation. Compared with traditional saliency maps, prototypical examples, and concept attributions, clinicians reported that the synergistic presentation “help[ed] them better understand what the algorithm is looking for… it is self-explanatory that staining was the most important factor distinguishing positive nuclei from those marked negative, with some identifying other important factors such as shape [and] size of nuclei.” Clinicians’ subjective ratings reinforced this feedback, where the synergistic explanation format received the highest median scores for clarity, relevance, trustworthiness, and practical utility.
While graphical synergy can improve explanation coherence, integrating multiple explanation types can raise concerns about fidelity to the underlying AI model. To synthesize transitional examples between prototypical positive and negative Ki67 nuclei instances, Evans et al. (2022) combined data attribution methods with image morphing techniques (Gulshad and Smeulders, 2021). Because these synthetic interpolations are not grounded in training data, they raise faithfulness concerns that highlight the need for new XAI techniques that seamlessly integrate explanation types while preserving fidelity. Developing such techniques would support richer graphical synergies, such as combining example-based explanations with saliency highlights (Kenny and Keane, 2021) or integrating salience maps with human-interpretable concepts (Yang et al., 2024).
5.1.4. Explanations as Recommenders
Traditionally, XAI work focused on explaining how systems work by helping users to assess, replicate, or predict AI decisions. Emerging work now shifts toward explaining why AI systems merit use, especially by supporting users in solving real-world problems beyond simple classification. In these contexts, explanations act as recommenders, helping users understand what actions to take and why those actions are beneficial.
For example, Wambsganss et al. (2020) developed an adaptive AI-assisted learning system to assist students in writing persuasive arguments. To facilitate revisions, the system visualized argumentative structure, evaluated writing quality, and provided natural language feedback alongside highlighted areas for improvement. Students expressed a significantly greater preference for using the new system than traditional reference-based writing tools. Nilforoshan and Wu (2018) implemented a similar application using feature-based explanations to improve reviewer responses on social media.
Besides writing, algorithmic recourse represents another practical use case of explanations. For instance, Kim et al. (2024) developed an XAI interface that combined counterfactual examples with recourse plans to help dysarthria patients correct mispronunciations. Similarly, Wang et al. (2023b) introduced GamCoach, which provided loan applicants with counterfactual explanations recommending financial changes to improve loan approval odds.
Findings from Wang et al. (2023b) highlight a key point that aligns with our framework (outlined in Section 3.3): effective use of an AI system does not necessarily require a deep understanding of system behaviors. For example, in their user study, participants rated “ease of understanding” (4.90) and “help with understanding the model” (5.17) lower than their “intention to reuse the system for future loan applications” (5.34) and “the system’s usefulness in identifying ways to improve their loan applications” (5.61), based on a 7-point Likert scale. Although observed differences were not statistically validated, this finding suggests that a deep understanding of model behaviors is not a prerequisite for effectively using the AI for algorithmic recourse. This reinforces our position on XAI’s rhetorical role: explanations should reveal how AI works and why an AI merits use for practical utility.
5.1.5. Explanations as Retrievers
Explanations can also function as retrievers to support dataset exploration and analogical reasoning. Different types of explanations highlight distinct model signals (see Appendix A), including local feature attributions, global patterns, representative prototypes, and human-interpret concepts. Users can leverage these characteristics to retrieve analogous cases, facilitate structured comparisons along shared attributes, and derive actionable insights grounded in prior experience. Building on this idea, researchers have developed interactive case-based retrieval systems that surfaced prior instances based on similar features or concepts through explanations. Applications include clinical diagnosis review (Cai et al., 2019b; Suresh et al., 2022), auditing dataset annotation quality (Kenny et al., 2021a), identifying model vulnerabilities (Sun et al., 2023a), and refining chatbot behaviors (Miranda et al., 2021).
Consider Cai et al. (2019b)’s case-based retrieval tool for medical diagnosis. Rather than evaluating AI outputs directly, pathologists used the tool to refine search queries that retrieved similar scans with confirmed diagnoses, supporting their interpretation of the current image. Pathologists improved queries by either cropping salient regions from normative scans or adjusting the feature importance of known medical concepts. Here, the primary XAI objective is to improve diagnostic accuracy for pathologists (i.e., why AI merits use) rather than understanding and verifying AI predictions (i.e., how AI works).
5.2. Credibility
Explanations can influence users’ perceived trust in AI systems when designed to project credibility. For the first goal of understanding how AI works, we identify three credibility-focused design strategies: Authority Framing, Cultural Relatedness, and Numerical Association. To communicate why an AI system is trustworthy and worth adopting, we identify two additional strategies: Technical Assurance and Social Endorsement.
5.2.1. Authority Framing
HCI researchers have examined how different communication styles affect user acceptance of machine-generated advice. These styles include direct versus implicit recommendations (Rau et al., 2009) and languages that signal authority or expertise (Metzger et al., 2024; Saunderson and Nejat, 2021). Naturally, designers have integrated authoritative language and tone into explanations to enhance their perceived credibility (Cambria et al., 2023b; Okoso et al., 2025). We refer to this design strategy as Authority Framing.
For example, an AI-recommended insulin dose can be phrased persuasively to make patients more likely to trust and act on the recommendation. In a detached or impersonal tone, the technical explanation might read:
“Because your planned alcohol intake exceeds one unit, the system recommends a lower dose of insulin. A normal dose would have been appropriate if the intake were one unit or less.”
(Example adapted from van der Waa et al. (2021))
Alternatively, the same explanation can be delivered in a more supportive tone, echoing a nursing communication style that balances empathy with expertise.
“Planning to enjoy more than one drink? The system suggests making a small adjustment to your insulin dose to keep you feeling your best. If you only have one unit or less, your usual dose is just right. We’re here to support you in making healthier choices!”
Another example is an AI tour guide by Díaz-Rodríguez and Pisoni (2020), which combined expressive language with domain expertise to explain the historical context of paintings, sculptures, and other artworks. Visitors responded more positively when explanations conveyed both factual accuracy and emotional resonance.
5.2.2. Cultural Relatedness
Beyond linguistic styles, XAI researchers have also integrated visual elements (e.g., color and shape) into explanations to project a credible tone, voice, or personality. Users are more likely to understand and trust explanations when presented visual elements resonate with their beliefs, values, or cultural identities. We refer to this design strategy as Cultural Relatedness.
For example, Okolo et al. (2024) studied the deployment of a neonatal jaundice [^3] Neonatal jaundice is a prevalent health condition in newborns, marked by yellowing of the skin and eyes. detection application in rural Indian clinics. During a co-design workshop, community health workers raised concerns about the default color scheme used in LIME (Ribeiro et al., 2016) and SHAP (Ghorbani and Zou, 2019) visualizations. In these explanatory visualizations, red simply denoted a classification outcome. However, health workers found the red color schemes confusing because red typically signifies severity and danger in medical contexts. To address this concern, workers applied traditional Hindu colors and styles to replace default color schemes in LIME and SHAP.
To avoid misinterpretations associated with colors, Okolo et al. (2024) experimented with other visual techniques to convey feature importance, such as textual descriptions and different shapes (e.g., squares and circles). Clinicians found that replacing color with text improved clarity, but using shapes caused new confusion, as some interpreted quantity (e.g., number of circles) as a proxy for condition severity. These findings emphasize the need for culturally informed visual designs to avoid unintended miscommunications.
5.2.3. Numerical Association
Another common technique to project credibility is numerical association, where explanations embed numbers, scores, or other metrics to reinforce perceived objectivity and authority. Numbers often carry a rhetorical power: they signal precision, rigor, and algorithmic sophistication. Users tend to equate numerical information with higher system competence, even if they do not fully understand its statistical meaning (Li et al., 2023).
For example, Ehsan et al. (2024a) implemented three different styles of AI explanations to evaluate how users perceive an AI agent’s intelligence. Regardless of participants’ backgrounds in AI knowledge, users consistently rated the agent with numerical explanations as more intelligent and reliable, even when they lacked a clear understanding of the presented metrics. The authors attributed this unwarranted faith to preconceived associations between numerical information and algorithmic thinking processes, reinforcing a perception of objectivity without true interpretability.
These examples highlight a key challenge: managing the tradeoff between credibility (ethos) and logical clarity (logos). While design strategies like cultural relatedness and numerical association can increase users’ perceived trust in AI systems, they do not always support a better understanding of how the system works. In some cases, as observed in Okolo et al. (2024) and Ehsan et al. (2024a), credibility-focused strategies may even obscure how AI systems work by introducing misinterpretations or inducing false confidence.
5.2.4. Technical Assurance
In practice, AI developers often employ technical criteria to validate predictive accuracy and fairness. Examples of technical standards include mitigating model bias (Gupta et al., 2023) and ensuring robustness against adversarial inputs (Gupta et al., 2024). Incorporating these criteria into explanations can foster perceived user trust by clarifying the AI system’s legitimacy and justifying why it merits use. For instance, Mitchell et al. (2019)’s model cards and Ribeiro et al. (2020)’s model behavior checklists exemplify how technical reliability can be assessed and communicated to users. Their design strategies convey an implicit message of assurance: We have made every effort to ensure this AI model is reliable so that you can use it safely and confidently.
While primarily for developers, tools like model cards and behavior checklists also serve a broader purpose. They signal that the model aligns with rigorous technical standards, which can subtly influence perceived user trust and adoption in practice. Like car manuals, which are rarely used but still inspire confidence, the real value of these technical artifacts lies in conveying assurance and reliability. Moreover, these artifacts also help brand a company’s identity around commitments to quality and responsible innovation.
To support complementary human-AI teaming, designers have incorporated technical criteria alongside adaptive explanations (introduced in Section 5.1). For example, Zhang et al. (2020) incorporated AI confidence scores alongside feature-based explanations to examine whether displaying confidence levels would increase users’ likelihood of accepting AI predictions. Similarly, Bansal et al. (2020) adapted the presentation of explanations based on the system’s confidence, showing explanations only when the model’s predicted confidence was high. Both studies found that participants heavily relied on these technical signals to guide their acceptance of AI predictions.
However, when exposed to technical indicators such as confidence scores, users may blindly follow AI predictions without critically assessing explanations. Despite this concern, researchers have yet to identify the most effective strategy to mitigate over-reliance. Some researchers have explored adding cognitive forcing functions to encourage more critical engagements with AI explanations (Gajos and Mamykina, 2022). Other researchers attribute the functional utility of explanations to domain experts, suggesting that their effectiveness depends on the user’s prior knowledge (Bansal et al., 2020; Wang and Yin, 2021). These works highlight a key trade-off between AI explanations’ functional utility and persuasive credibility. If users ultimately learn to perform the task independently, what added value does the AI system continue to provide?
We argue that engineers and technical developers should be responsible for ensuring that technical criteria support the responsible use of AI. Given the limited expertise of lay users, it is unrealistic to expect them to critically evaluate the reliability of technical signals. Once technical signals can faithfully convey the AI’s reliability, they should be carefully integrated into user-facing designs to promote assurance and confident use.
5.2.5. Social Endorsement
Beyond incorporating technical assurances, researchers have also employed social endorsements to strengthen users’ perceived trust in AI systems. In some cases, social affirmations exert a greater influence than the explanations themselves on whether users accept or reject AI decisions.
For example, Ehsan et al. (2021) developed an explainable AI interface for sales recommendations. Alongside feature-based explanations, the system provided a historical trajectory that showed how past sales teammates responded to the AI recommendation and the rationales behind their choices. Participants reported that these design elements enhanced their sense of social validation, transitive trust, and the temporal relevance of AI recommendations within the organizational setting.
While Ehsan et al. (2021) framed their design as promoting sociotechnical transparency, we interpret their approach as an example of social endorsement. Similar to how product reviews influence consumer behavior (Talib and Saat, 2017), their design leverages peer adoption to foster trust. When new users see that trusted colleagues have adopted and endorsed the system, they are more inclined to perceive the AI as credible and worth using by following their peers.
5.3. Emotional Resonance
Emotion has historically played a crucial role in design (Ho and Siu, 2012; Norman, 2007), serving essential communicative and experiential functions. In XAI, researchers have incorporated affective elements into explanations to make them more engaging. We identify three design strategies to support explaining how AI works: Explanation Affect, Emotional Cues, and Protective Framing. Additionally, we identify two strategies to explain why AI applications are emotionally compelling to use: Expectation Management and Expressive Interaction.
5.3.1. Explanation Affect
Different types of AI explanations can inherently evoke varying emotional responses depending on their relevance to the task, a phenomenon known as explanation affect. For example, Bernardo and Seva (2023) implemented example-based, feature-based, and rule-based explanations for animal identification. Participants judged the emotional appeal of each explanation type, highlighting differences in how each form engaged users. The study reported that example-based explanations elicited stronger feelings of surprise, trust, and perceived usefulness. In contrast, feature-based and rule-based explanations led to reduced emotional engagement.
However, the authors cautioned that these findings might not generalize to other tasks. In their study, the red color used in feature-based explanations to mark the Tarsier’s eyes and ears might have unintentionally evoked negative emotions such as fear or discomfort, an effect absent in the presented example-based explanations. Additionally, the author emphasized that users’ prior emotional states and attitudes toward AI could strongly influence how different explanation formats affect decision-making and trust.
5.3.2. Emotional Cues
Unlike explanation affect, where different types of explanations inherently evoke diverse emotions, some XAI researchers have directly integrated emotional cues into explanations to enhance user engagement. Examples of emotional cues include emojis, color, and tone of language.
For instance, Freire et al. (2023) incorporated emojis into AI explanations to increase user engagement. Bernardo and Seva (2023) compared robotic versus human-like language styles and found that robotic phrasing heightened anxiety and reduced perceived usefulness, while human-like language eased fear and improved perceived utility. In some cases, however, emotional responses are triggered unintentionally. For example, Okolo et al. (2024) found that the red color used in LIME and SHAP visualizations evoked negative emotions in a medical context, where red is commonly associated with danger, even though this was not the intended message of the explanation.
Additionally, researchers have incorporated emotional cues into interactive AI explanations. For example, in Wang et al. (2023b)’s GamCoach (introduced in Section 5.1.4), the design leveraged different colors, emojis, and visual cues to recommend interactions for algorithmic recourse. Participants found the tool relatively easy and enjoyable to use.
5.3.3. Protective Framing
Emotional cues are typically designed to elicit positive reactions toward AI explanations, making them more engaging, approachable, or enjoyable. In contrast, protective framing [^4] We adopt this term from Fokkinga and Desmet (2012), who explored how people use negative emotions in everyday life to achieve goals relies on negative emotions like skepticism or caution to encourage more critical reflection of AI explanations. Our narrative literature review did not identify design examples that explicitly employed this strategy. Instead, many studies introduced algorithmic aversions or AI system failures as independent variables to examine user acceptance of AI decisions and engagement with explanations. Thus, we propose protective framing as a possible direction for future XAI design.
For example, Ebermann et al. (2023) and Ben David et al. (2021) found that negative feelings such as disappointment, annoyance, or frustration often arise when users encounter unexpected AI outcomes or failures. These adverse emotions can heighten users’ attention to explanations. Notably, Ben David et al. (2021) observed that following an AI failure, participants were significantly less likely to accept subsequent AI advice when explanations were shown, compared to when no explanations were shown. This suggests that explanations can amplify critical scrutiny under negative affect.
Building on this observation, we could imagine a design scenario illustrating how protective framing could facilitate more critical user engagement. When the AI prediction contrasts with a user’s expectations, a system alert might say:
“The AI has predicted a different outcome, but it’s important to remember that both humans and AI can sometimes make mistakes. We encourage you to review the detailed explanation behind the AI’s decision…[show AI explanations]”
Even when the AI prediction aligns with the user’s decision, the system could still deliver a cautionary alert, reminding users of the potential for AI to make errors and encouraging them to review the explanation critically. In this use case, adding cognitive forcing functions (i.e., to slow down the decision-making process) can enable users to process explanations more carefully (Gajos and Mamykina, 2022; Buçinca et al., 2021).
In AI-assisted decision-making, failures could serve as a protective frame, prompting users to explore AI explanations proactively before finalizing their decisions. This idea reflects an alternate interpretation of the cost-benefit model described by Vasconcelos et al. (2023). While their work focused on reducing the cost or increasing the benefits of verifying AI predictions (through explanations), triggering negative emotions emphasizes the consequences of neglecting verification, effectively increasing the perceived benefit of careful review.
5.3.4. Expectation Management
In real-world AI applications, explanations are commonly used to manage user expectations. Before interaction, explanations can help justify why the AI system is worth using by outlining its strengths and limitations under intended use cases. After interaction, explanations clarify how the AI works by outlining the factors influencing its recommendations. This mirrors how people use explanations to set and adjust expectations, maintain shared understanding, and resolve ambiguity in everyday communication.
For example, Kocielnik et al. (2019) found that presenting AI explanations upfront could lower user expectations of AI performance, thereby reducing dissatisfaction during AI breakdowns. However, this emotional remedy may not be effective when explanations are not perceived as practically helpful. For example, Ebermann et al. (2023) found that participant mood scores significantly declined when explanations justified alternative decisions rather than the AI’s predictions, creating cognitive dissonance and undermining user satisfaction. Interestingly, despite the observed negative emotional impact, participant evaluations of system usefulness (i.e., self-rating on the question “to what extent was my [AI] support helpful”) actually improved. This divergence suggests that emotional discomfort does not always indicate poor system design. Like how students may resist challenging coursework yet still benefit from it, friction in human-AI interaction may play a productive role by encouraging more critical system evaluations.
5.3.5. Expressive Interaction
Expectation management primarily treats explanations as affective band-aids to remediate AI breakdowns. While this reactive use of explanations can be effective, designing with emotional elements also calls for a more proactive and holistic perspective. Beyond remediation, AI explanations can enable expressive interaction, a design approach that sustains emotional engagement across the user’s interactive journey with the AI system.
For example, Alvarado et al. (2022) developed a humanoid controller to interact with a movie recommendation algorithm. Users can accept or reject AI-suggested movies through various physical controls, such as pressing, holding, or shaking the controller. Notably, a long press at the controller’s center triggers the system to explain why a particular movie was recommended based on the user’s viewing history or personal preferences. This gesture provides a subtle way to address user curiosity about why a particular recommendation was made. More importantly, AI explanations are woven into a larger embodied interaction, functioning as a natural part of the experience rather than as isolated add-ons.
Expressive interaction treats emotion not as a remedy, but as a generative force that actively shapes how users engage with AI systems. Rather than smoothing over frictions, this design strategy aims to seamlessly integrate emotional responses as a natural part of human-AI interplay. Through repeated interactions, users develop personal rhythms with the system: deciding when to seek explanations, how much to rely on them, and what emotions to attach to different types of feedback. In this way, explanations become part of an evolving emotional narrative, rather than static tools for understanding model behavior.
6. Discussion
By synthesizing design strategies employed in prior XAI systems, we have identified a pluralistic design landscape, reflecting how researchers have explained how AI works and why AI merits use. Since the reviewed literature captured only a subset of existing design strategies, our synthesis should be viewed as a starting point rather than a comprehensive catalog. As XAI evolves, new strategies will likely emerge for different rhetorical appeals. In this section, we discuss the contributions of these diverse strategies to existing XAI research (Section 6.1 and 6.2), the caveats of identified design strategies, and limitations of our review process (Section 6.3).
6.1. Centering XAI Around User Intent and Task Relevance
As described in our rhetorical framework (Section 3.3), we extend the traditional goals of XAI beyond merely clarifying model behavior to include another dimension: explaining why AI merits use. Note that this goal is better understood as a design objective rather than a technical optimization. As highlighted in Section 5, prior researchers have already employed various design strategies to support this objective. Some used explanations as recommenders or retrievers, offering actionable insights for real-world problem solving. Others applied credible or emotional framing to make AI applications more compelling and persuasive.
A natural direction for future XAI research is to incorporate user intent and task relevance more explicitly as design criteria when selecting and evaluating explanations. In particular, researchers might consider whether a user’s strong motivation to adopt AI for a given task creates a meaningful need for explanations, and to what extent that need varies by task contexts. Establishing this connection between user intent and explanatory relevance would, in turn, enhance both the interpretive value and empirical rigor of XAI research.
For example, when examining the effects of different AI explanations on sentiment classification, recidivism prediction, or other tasks, the first set of questions could start from: Who are the targeted users of these AI tasks, and why would they want AI to assist them? Answers to these questions could provide specific use cases to further investigate users’ explanatory needs: Why would they care about understanding individual predictions or model global behavior in these tasks? What AI explanations and design strategies could better achieve their goals for using AI? Even when traditional AI explanations fall short, answering these questions can reveal new user needs and prompt designers to rethink the problem space, leading to more relevant solutions (Zimmerman et al., 2022). This shift in perspective opens opportunities to repurpose XAI techniques to support emerging AI applications beyond their original intent (Liu et al., 2024).
6.2. Perceiving Explanation Design as a Reflective Rhetorical Practice
In HCI, there is growing interest in promoting stakeholder participation and integrating cross-disciplinary knowledge when developing XAI systems (Malizia and Paternò, 2023; Mohseni et al., 2021; Dhanorkar et al., 2021). Within these collaborative settings, designers can play a critical role in addressing diverse stakeholder needs by drawing on the strategies synthesized in Section 5. Specifically, designers could balance different rhetorical appeals to accommodate the diversity and flexibility of human needs. For example, in some use cases, logical comprehension of the AI system and its practical utility take precedence, with credibility and emotional factors serving a secondary role. This is often the case when designing AI-powered analytical tools for technical users, such as developers or data scientists. In contrast, for consumer-facing AI applications, credibility and emotional resonance may be more influential in enhancing user engagement and trust.
Additionally, different design strategies can help mitigate one another’s limitations. For example, Gaole et al. (2025) observed that users tended to overrely on the AI system when presented with conversational or mechanistic explanations, which created an illusion of explanatory depth. To address this, designers may incorporate emotional strategies such as protective framing or expectation management to encourage more critical user engagement.
However, integrating and balancing multiple design strategies to meet diverse stakeholder goals remains a complex, evolving challenge. For example, an XAI design solution that worked well at one point, or for a particular stakeholder group in a specific domain, may become inadequate as the application context evolves, necessitating new design responses (Friedman, 2008; Zimmerman et al., 2022). Buchanan (1992) argued that design is a form of inquiry, where the actual problems emerge through continuous user engagement, reflection, and iterative assessment. This perspective is particularly relevant for designing AI explanations, which increasingly involves collaborative and sociotechnical considerations (discussed in Sections 2.3 and 2.4). Thus, addressing this challenge may require designers to use an iterative, exploratory design process (e.g., research through design (Zimmerman and Forlizzi, 2014)) combined with creative strategies to navigate potentially competing and shifting needs.
6.3. Rhetorical Design Caveats and Limitations of Our Review Process
6.3.1. Misuses of Design Strategies
Some design strategies introduced in Section 5, such as using numerical associations to project perceived AI intelligence (Ehsan et al., 2024a), demonstrate that explanations can be intentionally manipulated to increase user adoption. For example, designers could disseminate unreliable signals suggesting strong social trust when user trust has not been fully established. This might involve falsely claiming that the system has undergone rigorous testing when the AI has not, thereby manipulating user trust under inaccurate pretenses.
This risk of manipulation recalls the analogy of designers as rhetoricians (introduced in Section 3.3), a role increasingly associated with negative connotations in marketing (Miles and Nilsson, 2018), adversarial propaganda campaigns (Farrell and Schneier, 2018), and foreign influence operations (Starbird et al., 2019). However, this is not our intent. Like Buchanan (Buchanan, 1985a, b), we position rhetoric as a means to support reasoned invention and explore meaningful design alternatives. Future research can explore balancing rhetorical design strategies with the ethical concerns of potentially misleading target audiences. One approach could involve identifying misguided designs based on the three rhetorical appeals (e.g., logos, ethos, pathos), framing them as dark patterns (Gray et al., 2018) to warn future XAI designers.
Additionally, by reconsidering the role of rhetoric in explanation design, we argue that design strategies should not be viewed as inherently deceptive but can instead support human-centered values and responsible use of AI. Designers should intentionally ground rhetorical explanations in principles such as sustainability (Blevis, 2007), dignity (Kim, 2021; Buchanan, 2001b), and social inclusiveness (Costanza-Chock, 2020; Thorpe and Gamman, 2011), thereby mitigating the risk of misuse and reinforcing ethical engagement with AI systems.
6.3.2. Limitations of Our Review Process
Future work can refine our literature review process as rhetorical design gains broader recognition in the XAI community and new design strategies are proposed and systematically evaluated.
In our review, we categorized papers according to two primary XAI goals: (1) explaining how AI works, and (2) explaining why AI merits use. We observed that many studies targeting the first goal also included measurements relevant to the second, such as perceived AI usefulness and user trust in the AI system. For example, technical assurance was frequently incorporated in XAI studies to measure how users perceived different explanations, even though these studies did not explicitly articulate user intent and the task relevance (e.g., why participants might use AI to complete the task). Nonetheless, our axial coding incorporated both quantitative and qualitative findings from prior studies to understand the influence of different design strategies in promoting the AI’s benefits. Future reviews might consider more stringent inclusion criteria to ensure more precise alignment between study design, user intent, and task relevance.
Additionally, as described in our axial coding process (Section 4), we mapped user study measurements from prior work to the three rhetorical appeals in our framework (Section 3.3). However, this mapping process was inherently conceptual and subjective, incorporating author biases. On the other hand, this mapping also suggests an opportunity to develop a comprehensive set of measurement tools explicitly aligned with the dimensions of logical reasoning, credibility, and emotional resonance. Future research can systematically examine how existing evaluation metrics align with each rhetorical appeal, followed by a scoping review to understand the effects of different design strategies better.
7. Conclusion
In this paper, we propose conceptualizing AI explanations as rhetorical artifacts, designed not only to support user understanding of AI systems but also to promote appropriate persuasive adoption as another dimension of responsible AI. Drawing on the concept of rhetorical design, we develop a rhetorical framework that guides our coding process for a narrative literature review. Through this review, we synthesize a range of design strategies used in prior XAI studies to make explanations logically convincing, credibly appealing, and emotionally engaging. By critically examining the strengths and limitations of these strategies, this paper extends the design space of XAI by providing designers with new tools to address the complex needs of diverse stakeholders and the broader sociotechnical challenges involved in building XAI systems.
Acknowledgments
We thank Ahmer Arif, Jacek Gwizdka, James Howison, Yujin Choi, and Yiwei Wu for their valuable comments. This research was supported in part by Good Systems [^5] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsors.
Appendix A Types of AI Explanations
We briefly summarize AI explanations designed in a static form, including example-based, feature-based, counterfactual, rule-based, and concept-based explanations. These explanations serve as the basic design unit for more complex and interactive AI explanations (Bertrand et al., 2023). Note that this is a design taxonomy based on a logical understanding of how AI works, which might not closely align with the taxonomy used by AI developers, who might prefer the classification of XAI based on interpretable techniques (see Section 2.1).
A.1. Example-based Explanations
Example-based explanations generally refer to the XAI methods that use specific instances that are either retrieved from the training dataset or generated as new data points to explain how AI makes a particular decision. As Molnar (2020) and Wang et al. (2019) emphasize, the benefit of these explanations lies in how humans often use examples or analogies to aid in decision-making.
Researchers have explored different kinds of examples to explain how AI works. The most commonly selected examples are those that share similar features to the input. We refer to these examples as nearest-neighbors examples because “nearest neighbors” are both intuitive and widely used as a common technique for researchers to retrieve such examples. The classes of these nearest-neighbors examples can vary to the predicted output class depending on whether nearest neighbors are directly used in the model influence function. For example, if the prediction model is a k-nearest neighbors model, the examples should share the same output class as the input (Papernot and McDaniel, 2018; Su et al., 2023). However, if other prediction models are used and nearest neighbor is only used as a retrieval technique, the output class of these examples might differ (van der Waa et al., 2021; Cai et al., 2019a; Yang et al., 2020; Wang and Yin, 2021; Chen et al., 2023b). Therefore, it’s important to distinguish these two use scenarios to communicate how AI works. Using nearest neighbors as a retriever can only be interpreted as using examples to examine AI predictions. It might be incorrect to claim that these examples directly influence AI predictions.
Cai et al. (2019a) describe another two kinds of examples based on the class difference of nearest-neighbors examples. They are comparative examples, examples similar to the input but with alternative output classes, and normative examples, prototypical examples with the same class to the input. Both yield a different reasoning process for humans to map AI inputs to outputs based on examples. The idea of comparative examples is similar to counterfactual explanations, informing us how an instance could be changed to alter the AI prediction (Molnar, 2020; Verma et al., 2024) (We elaborate this in Section A.3, counterfactual explanations). In contrast, a normative example serves as a benchmark, allowing people to compare the input with a prototypical example to understand how their similarities result in the same output class.
A.2. Feature-based Explanations
Feature-based explanations generally refer to showing feature attribution or influence by indicating which input feature is important to the model global behavior or how it has direct positive or negative influence towards different output classes (Molnar, 2020; Wang et al., 2019; Mohseni et al., 2021). Unlike example-based explanations that are mostly local interpretations (i.e., how a specific input affects the output based on examples), feature-based explanation can communicate either local or global interpretation (i.e., how AI behaves specifically in a particular instance or generally on all features) depending on using which XAI techniques (Molnar, 2020).
Local feature contribution, which is also named as feature salience, highlights input features showing their contribution for different output classes. In other words, it tells us how each feature directly affects a model to make a negative or positive prediction for the test instance (Ribeiro et al., 2016; Molnar, 2020). In contrast, global feature importance or named feature relevance, would rank input features based on their overall importance that influence the model behavior at large (Lundberg and Lee, 2017; Molnar, 2020). In other words, it indicates how the model relies on different features for making future predictions in general. These two types of feature-based explanations work in a complementary fashion to communicate how AI works based on feature attributions.
A.3. Counterfactual Explanations
Counterfactual explanations characterize a hypothetical scenario telling us what could have happened if something changes (Verma et al., 2024). By adapting this scenario to describe the model’s input-output relationship, counterfactual explanations help inform the smallest change to the input values that alter the model prediction (Verma et al., 2024). Molnar (2020) highlights that the presentation of counterfactual explanations are always example-based as they have to be different instances showing how to change the test instance.
Based on the above definition, the differences between counterfactual explanations and several types of example-based explanations described earlier could become much clearer. First, we can infer that comparative examples are always counterfactual but counterfactual explanations do not need to be actual examples existed in the training data. They can be hypothetical examples by changing feature values. More importantly, counterfactual explanations infer actionable changes, as Verma et al. (2024) said “[they] do not explicitly answer ‘why’ the model made a prediction; instead, they provide suggestions to achieve the desired outcome.” This suggests that the counterfactual effect of a comparative example is demonstrated by informing people what features have been changed, thereby altering model prediction. Only displaying a comparative example without highlighting the feature changes from the original example might not result in a perceivable and understandable counterfactual effect.
A.4. Rule-based Explanations
Rule-based explanations typically involve explaining model predictions by outlining one or several decision rules (Molnar, 2020). These rules might take the form of a simple “if…then…” statement or more complex logic using a combination of arithmetic operators, such as “AND…¿…=…”. These rules provide a clear and structured way for humans to understand how the model arrived at a specific decision.
Rule-based explanations can inform either local or global interpretations, depending on the specific techniques used to generate these rules (Aghaeipoor et al., 2023). For example, interpretable rule-based machine learning models, like decision rules, are trained on all instances to extract rule sets as model decision-making conditions (Molnar, 2020). In contrast, model-agnostic techniques like Anchors (Ribeiro et al., 2018) create rules that explain the relationship between inputs and outputs for instances close to a specific test instance. These nearby instances are created using a perturbation approach. By focusing on these neighboring instances, these rules only characterize how certain features of the test instance specifically affect the prediction, regardless of changes in the other feature values.
A.5. Concept-based Explanations
Concept-based explanations primarily present user-defined concepts embedded within the model’s latent space to explain its predictions in the form of natural language or a ranking of concept significance (Kim et al., 2018; Molnar, 2020). For example, a zebra (the image class predicted by AI) is generally identified by its black-and-white stripes (identified concepts).
Typically, concept-based explanations shed light on a model’s global behavior by recognizing and extracting shared semantic features for a particular class, such as colors or patterns, from the model embedding. More recently, local approaches are proposed, such as identifying the most influential regions of an image then associating these regions with broader, global concepts (Shukla et al., 2023) or using prompt-based image segmentation model to automatically detect concepts within an image (Sun et al., 2023b).
Appendix B Articles for Review and Corresponding Coding Scheme
B.1. Reviewed articles
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Alqaraawi et al. (2020) Isolated presentation Logical: Subjective metrics Frequency of feature mentioned
Credible: Subjective metrics Self-confidence
Alvarado et al. (2022) Isolated presentation; Expressive interaction Credible: Qualitative feedback Increased sense of user control and awareness
Emotional: Qualitative feedback Increased sense of user enjoyment
Bansal et al. (2019a) Adaptive presentation Logical: Objective metrics User decision-making error boundary across trials
Bansal et al. (2019b) Adaptive presentation Logical: Objective metrics User decision-making error boundary across trials
Bansal et al. (2020) Adaptive presentation Logical: Objective metrics Human-AI agreement; Subjective metrics Explanation is useful, AI assistant is useful; Qualitative themes Used AI as a starting references point; Double-checked after they made own decisions
Credible: Qualitative themes Explicitly acknowledged they used the confidence [score]; Was more likely to accept AI above the [confidence] threshold
Ben David et al. (2021) Isolated presentation; Protective framing Logical: Objective metrics Acceptance of model advice
Credible: Subjective metrics Willingness to pay
Emotional: Independent variables User algorithmic aversion
Bernardo and Seva (2023) Isolated presentation; Explanation affect; Emotional cues Logical: Subjective metrics Perceived usefulness;
Credible: Subjective metrics Perceived trust, Reliance;
Emotional: Subjective metrics AI anxiety, Incidental emotion
Bove et al. (2022) Mechanistic presentation Logical: Subjective metrics User understanding of the system (e.g., local feature understanding, difference between local and global explanations)
Credible: Subjective metrics Explanation satisfaction
Cai et al. (2019a) Isolated presentation Logical: Subjective metrics System understanding, capability, and benevolence
Cai et al. (2019b) Mechanistic presentation; Explanation as retriever Logical: Subjective metrics Diagnostic utility, Future use, Overall preference; Qualitative themes Tracking the likelihood of a decision hypothesis, Generating new ideas; Reducing problem complexity
Credible: Subjective metrics Trust
B.2. Reviewed articles (continued)
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Cau and Spano (2025) Mechanistic presentation; Technical assurance Logical: Objective metrics Appropriate reliance; Subjective metrics User understanding and usefulness of explanations Qualitative themes Local explanations are fast and time-saving for decisions; Difficulty in understanding AI assigned weights on attributes, Overwhelmed by information
Credible: Subjective metrics User confidence, Trustworthiness of explanations; Qualitative themes Data-centric explanations were trustworthy with clear guidelines
Chen et al. (2023b) Isolated presentation Logical: Objective metrics Appropriate reliance; Subjective metrics Self-preferences; Qualitative themes Example-based explanations provide additional context and support inductive reasoning compared to feature-based; Example-based explanations led to less disruption of people intuition and provide strong signals of AI unreliability
Díaz-Rodríguez and Pisoni (2020) Authority framing; Cultural relatedness Credible: Qualitative themes Embedding expert knowledge for explanation
Emotional: Qualitative themes Communicating emotional depth in explanations
Ebermann et al. (2023) Isolated presentation; Protective framing Logical: Subjective metrics Evaluation of AI support
Credible Subjective metrics: Willingness to use AI
Emotional Subjective metrics: Mood; Independent variables: Cognitive misfit
Ehsan et al. (2021) Isolated presentation; Social endorsement Credible: Qualitative themes Increased social validation, transitive trust, and temporal relevance
Ehsan et al. (2024a) Isolated presentation; Numerical association Logical: Subjective metrics Perceived intelligence, Understandability
Credible: Subjective metrics Confidence, Friendliness, Second chance
Evans et al. (2022) Isolated presentation; Graphical synergy Logical: Subjective metrics Explanations are understandable, Explanations are useful; Qualitative themes Counterfactual explanations enable correct understanding of model behavior, Concept explanations seen as intuitive understandable but lack of precision; Saliency maps are visually appealing but risk promoting positive confirmation bias and obscure information.
Credible: Qualitative themes Confidence labels are intuitive cues for model reasoning
B.3. Reviewed articles (continued)
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Freire et al. (2023) Isolated presentation; Emotional cues Credible: Objective metrics Acceptance of AI advice
Emotional: Subjective metrics User engagement
Gaole et al. (2025) Adaptive presentation; Mechanistic presentation Logical: Objective metrics Objective feature understanding, Appropriate reliance based on human-AI agreement; Subjective metrics Perceived explanation utility
Credible: Subjective metrics User trust of the AI system, User confidence
Emotional: Subjective metrics: User engagement
Kenny et al. (2021b) Isolated presentation; Explanation as retriever Logical: Subjective metrics Perceived correctness, Reasonability
Emotional: Subjective metrics Trust, Satisfaction
Kim et al. (2024) Mechanistic presentation; Explanation as recommender Logical: Qualitative themes Increased usability for clinicians, Stronger desire to identify the types and causes of errors made by patients
Kocielnik et al. (2019) Isolated presentation; Expectation management Logical: Subjective metrics User understanding of AI model, Perceived control over system behaviors
Emotional: Subjective metrics Perceived system accuracy based on prior expectations
Liu et al. (2021) Isolated presentation Logical: Objective metrics Appropriate trust; Subjective metrics Real-time assistant useful, Training useful
Ma et al. (2023) Adaptive presentation; Technical assurance Logical: Human-AI agreement; Subjective metrics Perceived complexity, Future use, Perceived usefulness, Perceived helpfulness, Acceptable; Qualitative themes Participants in adaptive workflow were forced to think independently
Credible: Subjective metrics Trust in AI, Confidence, Explanation satisfaction, Trust in Correctness Likelihood; Qualitative themes Some participants distrust and ignore correctness likelihood but most participants referred to or were influenced by correctness likelihood based adaptation.
Miranda et al. (2021) Isolated presentation; Explanation as retriever Logical: Subjective metrics Model behavior understanding, Tool usefulness, Usability.
Nguyen et al. (2018) Adaptive presentation Logical: Qualitative themes Difficult to understand;
Emotional: Qualitative themes Enjoy playing with the tool
B.4. Reviewed articles (continued)
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Okolo et al. (2024) Cultural relatedness; Emotional cues Logical: Qualitative themes Color confused the interpretation of feature importance;
Credible: Qualitative themes Misinterpretation of patient severity due to default feature colors that conflict with medical domain norms;
Emotional: Qualitative themes Negative emotional perception
Okoso et al. (2025) Authority framing; Emotional cues Credible: Subjective metrics Acceptance of AI advice, Qualitative themes Some participants said tome influence their decisions and informal tone was inappropriate for a serious topic;
Emotional: Qualitative themes Some participants expressed discomfort with AI mimicking human emotions
Schemmer et al. (2023) Isolated presentation Logical: Objective metrics Appropriate reliance
Schuff et al. (2022) Isolated presentation Logical: Subjective metrics Self-reported importance ratings of features
Sun et al. (2023a) Isolated presentation; Explanation as retriever Logical: Subjective metrics Model understanding, Tool usability; Qualitative themes Insufficient description about reasonability matrix, Ill-formed local explanation visualization, Attention adjustment can be subjective
Suresh et al. (2022) Isolated presentation; Explanation as retriever Logical: Objective metrics Human-AI agreement; Qualitative themes Nearest neighbors enable reasoning with clinically-relevant concepts and help characterize model uncertainty, Feature-based explanations are difficult to extract high-level, clinically-relevant concepts;
Credible: Qualitative themes Visualizing variation helps assess prediction reliability
van der Waa et al. (2021) Isolated presentation Logical: Objective metrics Advice prediction, Decision factor identification; Subjective metrics Perceived system understanding, Perceived prediction of correctness
Vasconcelos et al. (2023) Isolated presentation Logical: Objective metrics Appropriate reliance, Task difficulty;
Credible (authors’ post-hoc reflections): Stakes of the task;
Emotional (authors’ post-hoc reflections): Enjoyability
Wambsganss et al. (2020) Adaptive & Mechanistic presentation; Explanation as recommender Logical: Subjective metrics Intention to use, Perceived usefulness, Perceived ease of use; Qualitative themes Positive feeling about adaptive learning given its fast and direct feedback, graph-like visualization, and score. However, sometimes the system is less accurate in claim and premise classification.
B.5. Reviewed articles (continued)
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Wang et al. (2019) Mechanistic presentation Logical: Qualitative themes Counterfactual reasoning is useful;
Credible: Qualitative themes Participants felt less trustworthy without the patient context presented in the explanations
Wang and Yin (2021) Isolated presentation Logical: Objective metrics Human understanding of model behaviors, Appropriate trust; Subjective metrics Self-reported understanding of model behaviors
Wang and Yin (2023) Adaptive presentation Logical: Objective metrics Frequency of accepting model decisions; Subjective metrics Perceived accuracy;
Credible: Subjective metrics Self-reported trust gain, Satisfaction gain
Wang et al. (2023b) Mechanistic presentation; Explanation as recommender; Emotional cues Logical: Subjective metrics Easy to use, Easy to understand, Help understand model, Help find ways to improve plans, Will apply loan again, Usefulness of specific interactions; Qualitative themes Controllable, Requiring small changes or less compromise, Beneficial for life;
Emotional: Subjective metrics Enjoyable to use
Yang et al. (2020) Isolated presentation Logical: Objective metrics Appropriate trust; Subjective metrics Explanation helpfulness, Usability;
Credible: Subjective metrics Self-confidence, Trust meter before and after AI feedback
Zhang et al. (2020) Isolated presentation; Technical assurance Logical: Objective metrics Human-AI agreement