### v2

# Rhetorical XAI: Explaining AI’s Benefits as well as its Use via Rhetorical Design

Abstract.
We explore potential benefits of incorporating Rhetorical Design into the design of Explainable Artificial Intelligence (XAI) systems. While XAI is traditionally framed around explaining individual predictions or overall system behavior, explanations also function as a form of argumentation, shaping how users evaluate system perceived usefulness, credibility, and foster appropriate trust. Rhetorical Design offers a useful framework to analyze the communicative role of explanations between AI systems, and end-users, focusing on: (1) logical reasoning conveyed through different types of explanations, (2) credibility projected by the system and its developers, and (3) emotional resonance elicited in users. Together, these rhetorical appeals help us understand how explanations influence people perceptions and facilitate AI adoption across and within different collaborative and social contexts. This paper synthesizes design strategies from prior XAI work that align with these three rhetorical appeals and highlights both opportunities and challenges of integrating rhetorical design into XAI design.
Explainable AI, Rhetorical Design
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing HCI theory, concepts and models
1. Introduction
Modern AI systems are notoriously opaque, limiting efforts to understand or audit their behaviors (Cambria et al., 2023a; Saeed and Omlin, 2023). In response, Explainable Artificial Intelligence (XAI) aims to foster trust and accountability in AI systems by making their decision-making processes more transparent (Busuioc, 2021; Jacovi et al., 2021). XAI encompasses techniques to support human understanding of both local predictions (individual decisions) and global behaviors (Mohseni et al., 2021; Cambria et al., 2023b; Cabitza et al., 2023; Lai et al., 2023).
Different stakeholders leverage explanations to pursue distinct goals (Kim et al., 2024; Hoffman et al., 2021; Preece et al., 2018; Tomsett et al., 2018; Mohseni et al., 2021). For example, developers employ explanations to improve predictive performance by debugging model errors and biases (Hohman et al., 2018; Molnar, 2020). Ethicists and policymakers rely on explanations to assess the safety and accountability of AI systems (Busuioc, 2021). HCI researchers investigate how explanations can foster user appropriate trust (Zhang et al., 2020; Lai et al., 2023; Vasconcelos et al., 2023) and support complementary human-AI teaming (Bansal et al., 2021). Stakeholders with diverse objectives tend to evaluate explanations through different lenses, shaped by their specific roles and needs. In particular, end-users tend to judge explanations alongside more pragmatic and affective dimensions such as utility, trustworthiness, and alignment with personal values (Vera Liao et al., 2022). Therefore, a faithful and logically accurate explanation that neglects human-centered considerations may be completely disregarded.
In this paper, we argue that designing explanations for effective adoption motivates expanding the scope of XAI beyond explaining how AI works towards also articulating why AI merits use. An AI system is just one of many possible solutions to a user problem and AI explanations can serve a rhetorical function by communicating why such a system is beneficial, credible, and appropriate in a given context. Applications might include supporting critical evaluations of AI decisions (Wang and Yin, 2022; Bansal et al., 2021; Nguyen et al., 2018), assessing system credibility (Binns et al., 2018; Dodge et al., 2019), and calibrating appropriate user trust (Markus et al., 2021; Liao and Sundar, 2022).
Toward this end, we bridge Rhetorical Design with XAI. Rhetorical design provides an analytical framework to analyze both the how and the why of explanatory goals through three dimensions: (1) logical reasoning, how different explanation types appeal to distinct forms of reasoning; (2) credibility, how the AI system and its developers convey trustworthiness; and (3) emotional resonance, how different affective reactions elicited by explanations influence user perception and engagement with the system. Guided by this framework, we use a deductive-inductive approach to select, code, and analyze design examples from prior XAI research. Through this process, we synthesize design strategies across the three rhetorical dimensions. These identified strategies are situated in two explanatory goals: (1) How AI works, and (2) Why AI merits use.
Our paper offers both conceptual and practical contributions to the design of XAI systems. Conceptually, we propose a rhetorical framework for XAI that expands the traditional focus on explaining how AI works to include why AI merits use. Practically, guided by this framework, we analyze existing design strategies through the three classical rhetorical appeals (logos, ethos, and pathos), highlighting how explanations can be designed to be more logically convincing, credibly appealing, and emotionally engaging. By integrating these communicative functions, we show how AI explanations can extend beyond technical transparency to become richer rhetorical artifacts that help diverse stakeholders interpret AI outputs, negotiate responsibilities, build trust, and coordinate decisions in collaborative and social contexts. By critically assessing the strengths and limitations of these strategies, we also discuss their broader implications for future XAI research and design.
This paper is organized as follows. In Section 2, we review key technical concepts and existing design paradigms in XAI, laying the foundation to highlight the missed opportunity of integrating rhetorical design into the current XAI landscape. In Section 3, we introduce and adapt rhetorical design into XAI through a conceptual framework. In Section 4, we describe how we apply our rhetorical framework to guide the qualitative coding process when reviewing prior XAI designs. We present our coding results in Section 5, elaborating on the identified design strategies. In Section 6, we critically examine the benefits and limitations of these strategies and offer recommendations for future XAI research. Finally, in Section 7, we summarize our contributions toward advancing rhetorical XAI.
2. Background
In this section, we briefly review key technical concepts (Section 2.1) and design paradigms (Section 2.2, 2.3, and 2.4) in XAI. This summary draws primarily from existing survey papers, position papers, and design frameworks in XAI, serving as background to contextualize our perspective on rhetorical design.
2.1. XAI Technical Concepts
Both AI and HCI researchers have acknowledged that making AI systems interpretable to people can foster more trustworthy and accountable use of AI (Markus et al., 2021; Molnar, 2020; Jacovi et al., 2021; Busuioc, 2021). Doran et al. (2017) categorize AI systems into three types — opaque, interpretable, and comprehensible — based on the level of transparency they offer in explaining automated decisions. In this taxonomy, interpretability refers to how well users can understand the internal processes by which an AI system transforms inputs into outputs. Other taxonomies of interpretability frame this input-output relationship through alternative lenses such as explanation scope, algorithmic techniques, or interactive format.
A common taxonomy based on the explanation scope is whether AI explanations are local or global (Molnar, 2020; Mohseni et al., 2021; Dwivedi et al., 2023). Local explanations explain why a specific input maps to an observed output. In contrast, global explanations provide an overarching view of how the AI generally maps inputs to outputs. Orthogonally, scope can also vary by model generality. Another foundational taxonomy differentiates model-specific techniques, which are tailored to a particular model, from model-agnostic techniques, which can be applied to any model.
Different XAI techniques lead to diverse types of explanations, which HCI researchers have evaluated with respect to their effects on users. These explanation types include example-based, feature-based, counterfactual, rule-based, and concept-based explanations (see Appendix A for further details). Each supports a distinct form of logical reasoning, reflecting the underlying algorithmic mechanisms that generate the explanation (Wang et al., 2019).
Given the wide range of human-AI interactions enabled by explanations (Bertrand et al., 2023; Xu et al., 2023; Chromik and Butz, 2021), many user studies adopt two primary forms of explanations — static vs. interactive — to examine how different explanation types influence user acceptance or foster appropriate trust in individual AI predictions. Static explanations typically present fixed outputs for given inputs and allow participants to engage with them over a limited number of trials. In contrast, interactive explanations enable users to actively engage with the system by adjusting inputs or model parameters, allowing them to explore alternative scenarios and gain a deeper understanding of the model’s behavior (Liu et al., 2021; Bertrand et al., 2023; Chromik and Butz, 2021; Nguyen et al., 2018).
While there is debate about the necessity of AI explanations, many researchers view XAI as a promising pathway for promoting safe, trustworthy, and accountable use of AI. For instance, when humans collaborate with AI to perform different tasks, explanations can help users understand model behavior, enabling them to judge when to accept or reject AI recommendations. This, in turn, fosters appropriate user trust and supports complementary human-AI teaming (Lee and See, 2004; Yang et al., 2020; Wang and Yin, 2021; Schemmer et al., 2023). To support cooperation, HCI researchers typically draw on human-centered principles when designing XAI systems (HCXAI). HCXAI emphasizes identifying user needs and contexts of use to inform whether, when, and how different explanations should be presented (Ehsan and Riedl, 2020; Ehsan et al., 2022b, c, 2024b).
Building on this perspective, the following sections briefly introduce XAI design paradigms including user-centered (Section 2.2), stakeholder-centered (Section 2.3), and sociotechnical approaches to explanation (Section 2.4).
2.2. User-centered Explanations
HCI scholars studying XAI have focused on how different types of explanations can support users in understanding AI systems more effectively (Hagras, 2018; Liao et al., 2020; Vera Liao and Varshney, 2021). Hagras (2018) first argued that XAI aimed to make AI understandable to humans. One strategy toward this goal is to reframe technical descriptions of AI decision-making into user-friendly language. For example, consider explaining how an autonomous car changes speeds. Instead of a technical explanation such as “if the distance to the car ahead is less than 2.5m and the road is 10% slippery then reduce car speed by 25%,” a more human-like natural language rephrase might be “if the distance to the car ahead is low and the road is slightly slippery, then slow down.” In general, researchers have developed various types of explanations (see Appendix A) that employ visualizations and natural language to make AI more accessible and understandable to people (Lai et al., 2023; Robbemond et al., 2022).
Given the diversity of explanations, researchers have proposed different frameworks to help designers select the most appropriate explanation type for a given context. For example, Liao et al. (2020) proposed an XAI “question bank” comprised of common user inquiries such as How AI works, Why AI made a specific prediction, Why not an alternative prediction, What AI would predict if the input changes, and How to change the input to alter the AI prediction. Other researchers have adopted this question-informed framework to select AI explanations when designing XAI systems (Xin He and Zhang, 2023; Ehsan et al., 2022a; Wang et al., 2023a). Other frameworks aim to bridge human understanding and machine explanations from a cognitive perspective, such as Wang et al. (2019) proposing to design AI explanations align with human cognitive processes, Chen et al. (2022) incorporating task-specific human intuitions when assessing user understanding of explanations, and more (Zhang et al., 2019; Subhash et al., 2022; Chen et al., 2023a).
Since AI systems serve diverse end-user groups, explanations can be tailored to their specific needs and contexts (Ehsan et al., 2024a; Kim et al., 2024). For example, loan seekers and lenders have different explanatory needs for understanding how a credit scoring AI system functions (Bove et al., 2022). Loan seekers might be particularly interested in why their financial record led to a rejection and how to alter the decision. Lenders might seek a global understanding of AI behavior across historical loan records to support financial analysis, enabling them to make more informed decisions about issuing loans to future applicants (Dikmen and Burns, 2022). Similar distinctions are also prevalent in the medical domain between patients and healthcare providers (Burgess et al., 2023; Kim et al., 2024). Collectively, these cases highlight how domain expertise shapes explanatory needs (Mohseni et al., 2021; Kim et al., 2024). Lay users often prefer personalized and actionable guidance, while experts commonly seek system-level understanding.
2.3. Stakeholder-centered Explanations
While end-users are a central focus in XAI, they represent just one of the many stakeholders involved in developing, deploying, and adopting AI systems. To situate AI within this broader socio-technical ecosystem, explanations should also cater to stakeholders’ distinct goals, responsibilities, and information needs. Two primary approaches for categorizing stakeholders are based on their roles within the AI lifecycle (Tomsett et al., 2018; Preece et al., 2018; Hoffman et al., 2021; Dhanorkar et al., 2021) and their level of AI knowledge (Mohseni et al., 2021).
In role-based stakeholder categorization, Hoffman et al. (2021) demonstrated distinct explanatory needs between developers, policymakers, and domain users. They reported that developers mainly employed explanations to debug AI errors and identify potential algorithmic biases to improve overall predictive performance. Policymakers leveraged explanations to audit AI behavior in expected deployment settings and unexpected edge cases. Domain users used explanations to understand or counterfactually remediate unfavorable AI decisions. Other researchers also used role-based frameworks but with alternative groupings of stakeholders (Tomsett et al., 2018; Preece et al., 2018; Dhanorkar et al., 2021).
Complementary to stakeholder roles, Mohseni et al. (2021) proposed a taxonomy of explanation needs based on varying levels of AI expertise. They distinguished between AI novices (layman users) from AI experts (engineers and developers), each leveraging explanations to suit different purposes. However, Suresh et al. (2021) contended that both role-based and AI knowledge-based frameworks could be problematic, as overlapping explanation needs could challenge consistency between different stakeholder groups. Consequently, they proposed to organize stakeholder needs around Goals, Objectives, and Tasks along the AI development cycle, aiming to offer a more cohesive and ontologically grounded vocabulary.
While stakeholder frameworks might differ in structure and granularity, they all highlighted the need to contextualize AI explanations to specific goals and responsibilities of different stakeholder groups (Wolf, 2019; Carroll, 2003). Identified stakeholder needs then served as specifications to guide the technical development (Kaplan et al., 2024) and interface design of XAI systems (Dhanorkar et al., 2021; Ehsan et al., 2023).
2.4. Socio-technical Explanations
As researchers broaden their study focus from end-users to a wider range of stakeholders, this shift reflects a socio-technical perspective for studying XAI (Preece et al., 2018; Ehsan and Riedl, 2020; Dhanorkar et al., 2021). Brandão et al. (2019) argued that AI experts often lack a natural understanding to connect the potential social meanings with models they create. Ehsan et al. (2023) further illustrated this disconnection through a use case where an AI-powered sales prediction tool, despite its 90% reported accuracy, was adopted by only 10% of sellers due to either over-trust by less experienced users or disregard by experienced sellers. Similar cases have emerged in different areas, including customer-service relationship (Northey et al., 2022), journalist norms and values (Komatsu et al., 2020), and frontline workers (Fox et al., 2023).
The main issue described in the above use cases is that, current AI deployment did not account for organizational social needs (Ehsan et al., 2023). In particular, users reported that verifying AI predictions often created added burdens for organizational work, such as having to learn new terminologies, consult technical experts, or manage concerns about reduced transparency. To address this, some XAI researchers argued that explanations should also account for users’ roles, expertise, and collaborative practices within the organization (Ehsan et al., 2021, 2023).
2.5. Research Focus of this Study
Thus far, Section 2 has outlined key technical concepts in XAI, covering commonly used terminologies and taxonomies (2.1), as well as design paradigms of human-centered XAI, including user-centered (2.2), stakeholder-centered (2.3), and sociotechnical perspectives (2.4). This section aims to differentiate our work by conceptualizing AI explanations as rhetorical artifacts that mediate communication between designers and end-users, beyond merely justifying outputs or revealing flaws.
While user-centered XAI (Section 2.2) prioritizes functional utility and cognitive human-AI alignment across different user objectives, we argue that pragmatic and affective responses equally shape understanding and adoption of AI (Dewey, 2005; Buchanan, 1985a, 2001a). For example, when interacting with XAI systems, users frequently grapple with questions such as: Will this explanation help me achieve my goal better? Can I trust its rationale? Does it align with my ethical principles of using AI? These questions highlight that technical fidelity and cognitive alignment alone are insufficient; explanations must also resonate with user expectations, emotions, and values to be meaningful and persuasive.
More importantly, while the stakeholder-centered framework (Section 2.3) broadens the scope of XAI to include the goals and values of broader stakeholder groups, and socio-technical focus (Section 2.4) highlights the practical challenges of implementing XAI in an organizational setting, both still frame AI explanations primarily as tools either to justify AI results or to expose their flaws. We argue that this framing neglects the role of explanation as a communicative act in AI use, which is shaped by design and situated within a social context, inviting diverse interpretations around why AI appears logically sound, credibly useful, and emotionally appealing. When multiple stakeholders engage with explanations in collaborative settings, it is important to understand how different explanation designs shape their interpretations. These interpretations, in turn, influence perceptions of AI transparency, the negotiation of responsibilities, the development of trust, and the coordination of final AI adoption in practice.
Our reframing positions AI explanations not only as informative responses but also as rhetorical artifacts that mediate the relationship between AI systems and end-users. In other words, AI is just one of many possible solutions to a user problem. AI explanations thus must serve a persuasive and contextualizing role by communicating why an AI system is beneficial, credible, and appropriate in different contexts. From this perspective, a rhetorical XAI design can more effectively support stakeholders in evaluating automated decisions (Wang and Yin, 2022; Bansal et al., 2021; Nguyen et al., 2018), assessing system credibility (Binns et al., 2018; Dodge et al., 2019), and calibrating appropriate user trust (Markus et al., 2021; Liao and Sundar, 2022).
In this paper, we investigate the argumentative role of AI explanations in mediating user adoption of AI systems, and we synthesize how prior research has employed various design strategies to fulfill this rhetorical function. We draw on the concept of Rhetorical Design, a longstanding product and graphic design practice that examines how form, structure, and presentation influence user adoption. In the next section, we introduce the theoretical foundations of rhetorical design in HCI and describe how we adapt this lens to examine and inform explanation design.
3. Rhetorical XAI
Originally developed in product design (Buchanan, 1985a), rhetorical design has expanded into graphic and communication design (Sosa-Tzec and Siegel, 2014; Sosa-Tzec et al., 2015; Foss, 2004; Barnard, 2013; Boyarski and Buchanan, 1994), information visualization (Hullman and Diakopoulos, 2011; Hua, 2021), and other HCI domains (Wahyurini, 2021; Rutsky, 1999; Brummett, 1999). Specific rhetorical strategies include visual metaphors and analogy, emphasis and contrast, information alignment and proximity, and more.
In this section, we first introduce the original concept of rhetorical design in product design (Section 3.1) and outline common strategies (Section 3.2) employed across design fields. Building on this foundation, we establish rhetorical design as an analytical framework (Section 3.3) for identifying design strategies from prior XAI research.
3.1. Rhetorical Design
Daniel Kahneman, Nobel Prize winner in Economics, has shown through his research that people often make decisions in ways that deviate from strict rationality (Kahneman, 2012). In product design, Buchanan (1985a) argued that product quality is not solely determined by functional aspects that reflected technical capabilities, but also by the credible signals and emotional cues embodied in its design. This is particularly evident when people choose a product not just for its utility, but also for factors such as its identity (Clark, 2009) or broader social significance (Stoekl, 2009; Forlizzi, 2008; Liu and Kim, 2023).
What makes a product compelling? To answer this question, Buchanan (1985a) reviewed a list of everyday products — from household utensils to electronics — then drew an analogy between a product and a verbal argument. He argued that a product functions as a rhetorical artifact designed to compel interest, attract attention, and invite use from its intended audience. From this view, a product’s rhetorical appeal extends beyond meeting practical needs by inviting users to envision new ways of acting, thinking, and being. Here, Buchanan drew on classical rhetoric (i.e., logos, ethos, and pathos) to contend that product design, like persuasive speech, seeks to influence how users act, think, and feel.
In classical rhetoric, Aristotle (1926) argued that an argument is persuasive not only due to logical reasoning (logos) but also because of the speaker’s credibility (ethos) and emotional appeal to the listener (pathos). For example, a genuine diamond ring (strong logos) may appear counterfeit if offered by a financially impoverished individual (weak ethos). Likewise, simple truths (strong logos) cannot be readily accepted by a child crying and throwing a tantrum (weak pathos).
Extending these rhetorical principles to product design, Buchanan (1985a) introduced Design Rhetoric, where logos refers to the product’s technological reasoning, ethos to the implied character or credibility of the manufacturer as reflected in the product, and pathos to the product’s usability and emotional appeal for intended users; Here, technological reasoning is the intelligent organization of materials and processes to solve practical problems in ways that appear logical, efficient, and appropriate to users’ needs and values. Effective designs balance these three rhetorical appeals through a designer’s own judgment (Buchanan, 2007).
To summarize rhetorical design, two main points emerge: 1) in product adoption, design functions as rhetorical arguments with specific strategies to make the product persuasive and compelling for use; 2) these strategies can be framed around three unique dimensions – logical reasoning, credibility, and emotional resonance – each aligning with the key actors in the rhetorical triangle: the product, the designer, and the user.
3.2. Design Strategies for Different Rhetorical Appeals
HCI practitioners have developed various strategies to fulfill the three rhetorical appeals. Here, we briefly review examples from User Interface (UI) design and data visualization, as they closely resonate with design practices in the XAI community. In web design, Ibrahim et al. (2013) observed that visually salient features such as “contrast, liking, or surprise” are effective in evoking emotional responses from users, credibility is supported by elements like “recognizability, assurance,” or “evidence and social proof” by showing peer activities on the web, and logical appeal is reinforced through features such as “simplicity, variety, and verifiability.” Similarly, Sosa-Tzec and Siegel (2014) analyzed how visual strategies and affective heuristics operate in graphical interfaces to elicit these rhetorical effects. They later published a rhetorical design handbook tailored for UI/UX contexts (Sosa-Tzec et al., 2015).
In information visualization, Hullman and Diakopoulos (2011) outlined a variety of linguistic (e.g., irony, similarity) and visual (e.g., visual metaphors, analogy, contrast, classification) strategies that enhanced narrative storytelling. They presented Mapping America [^1] Mapping America: Every City, Every Block. The New York Times: http://projects.nytimes.com/census/2010/explorer and Poll Dancing [^2] Poll Dancing: How Accurate are Poll Predictions? Datablog: https://thesocietypages.org/graphicsociology/2010/06/03/poll-data-poll-dancing/ as examples to illustrate how different strategies cohesively interact to make presented information more logically convincing, credible, and emotionally engaging. Many designers employed these strategies in infographics and visualizations to communicate complex social issues. For example, Hua (2021) used data association, embodiment, and visual metaphor to design infographics to raise public awareness about recycling contamination. Similarly, Meuschke et al. (2022) applied narrative-driven strategies to facilitate public understanding of medical data. Other directives include supporting civic engagement and participatory discourse (Dörk et al., 2013; Claes and Vande Moere, 2017).
This paper aims to synthesize design strategies (similar to those discussed above) for crafting AI explanations. In the following section, we introduce the rhetorical framework for identifying and categorizing these strategies, which guided our qualitative coding and review process of design examples from prior XAI research.
Dimensions Rhetorical Inquires for AI Explanations
Explain How AI Works Explain Why AI Merits Use
Logical
Reasoning
What forms of explanations are used, and how are they presented to convey the underlying logic of AI local prediction or overall behavior? How do these explanations help end-users apply AI effectively to address their real-world needs and tasks?
Credibility How does the framing and presentation of explanations signal their credibility to users? How do the framing and presentation of explanations signal the credibility of AI systems?
Emotional Resonance How do the framing and presentation of explanations engage user emotions to facilitate their understanding of AI? How do the framing and presentation of explanations engage user emotions to better utilize the AI system?
Table 1. A Rhetorical Inquiry Framework to Characterize AI Explanations.
3.3. A Rhetorical Framework to Characterize Explanation Design
Our framework builds on two primary ideas of rhetorical design, as outlined at the end of Section 3.1. Drawing on the first idea — in product adoption, design functions as rhetorical arguments with specific strategies to make the product persuasive and compelling for use — we propose extending XAI beyond merely (1) explaining How an AI system works (e.g., clarifying local predictions and global behaviors) to also (2) explaining Why an AI system merits use.
Our call to broaden XAI resonates with prior work, which argued that effective use of AI may not require a deep understanding of model behaviors. For instance, Miller (2021) noted, “Explanations of how models work might just distract people from figuring out what they really need.” Regardless of the specific needs (see Section 2.2), from a user-centered perspective, understanding how AI works (i.e., pure interpretability) is means rather than an end for why people choose to use AI in the first place. Therefore, explanations should go beyond the current focus on functional transparency, addressing more rhetorical questions such as: What is the compelling value proposition for a user to engage with AI?
Next, we apply the second idea — the rhetorical triangle of logical reasoning, credibility, and emotional resonance — as qualitative coding axes to guide our synthesis of design strategies in prior XAI literature. This framing also resonates with recent calls emphasizing that user perceptions of AI are influenced by more than just logic (Thommes et al., 2024). Experiential elements, such as tangible features (Ghajargar and Bardzell, 2023; Alvarado et al., 2022), embodied interaction (El-Zanfaly et al., 2023), and sensory or emotional responses of the AI system (Bryan-Kinns, 2024), can play a more significant role in influencing users to accept or reject AI recommendations.
Inspired by this line of work, we ask what if designers viewed themselves as rhetoricians? That is, by drawing on the rhetorical triangle: (1) logical reasoning conveyed through different types of explanations; (2) credibility projected by the AI and its developers; and (3) emotional resonance elicited in users, how might we craft compelling design arguments about AI through explanations? What design strategies exist, or need to be developed, to fulfill each rhetorical appeal and strike a thoughtful balance among them?
Building on this conceptual adaptation, we argue that rhetorical design can serve two complementary objectives in XAI: (1) supporting traditional AI explanations that clarify How AI Works, and (2) designing explanations that communicate Why AI Merits Use to guide or inform adoption. Our goal in foregrounding this distinction is not to enforce a rigid dichotomy, but to highlight the nuanced differences between these design objectives and propose corresponding strategies. More importantly, we hope that clarifying both objectives and keeping them in mind during the design process can enhance the value of explanations, suggesting which types of explanations to use and the contexts where they are appropriate or inappropriate regarding the technology affordance, task requirements, and the stakes involved (we discuss this further in Section 6.1).
We examine each objective through the three rhetorical dimensions: logical reasoning, credibility, and emotional resonance. In Table 1, we outline each rhetorical appeal as a coding inquiry across the two goals, which guided our open and axial coding process detailed in the next section. While Table 1 contrasts the two explanatory goals to clarify our coding process, this framing is intended to support, not discourage, explanation strategies that span both objectives.
4. Method
We conducted a narrative review guided by coding inquiries specified in Table 1. Unlike a systematic review, which follows established protocols such as PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) (Moher et al., 1996), we chose a narrative review to allow for greater flexibility in selecting, interpreting, and synthesizing diverse design strategies across different XAI contexts (Baumeister and Leary, 1997; Ferrari, 2015). We detail the review process in Appendix B, including search terms, databases, inclusion/exclusion criteria, and paper grouping strategies. Nevertheless, potential author bias remains an inherent limitation of this approach (see limitations in Section 6.3). To mitigate this, we employed a structured three-step process involving both open and axial coding.
Step 1 - Classifying Design Examples based on XAI Objectives: Before applying open coding to identify specific design strategies, we first conducted a top-down process to classify prior work based on whether AI explanations were primarily used to (1) convey how the AI system works or (2) enhance the usability and adoption of the AI. Specifically, we conducted a need-based meta-analysis grounded in the principles of user-centered XAI (Section 2).
For example, Cai et al. (2019a) presented example-based explanations designed to help users understand how an AI model recognizes a drawing of an avocado. Their primary motivation was to “help users develop intuition for the reasons underlying machine predictions, without directly exposing the internal logic of the algorithm.” Therefore, we classified Cai et al. (2019a) under the first XAI objective: explaining how AI works. In contrast, in another work by Cai et al. (2019b), the authors integrated example-based explanations and other interactive features into a case-based retrieval system to inform clinical diagnosis. We classified Cai et al. (2019b) under the second XAI objective since the presented explanations primarily demonstrate the practical benefits of the AI system. Additionally, we classified Wang et al. (2023b)’s interactive recourse tool for loan approval under both XAI goals, as explanations provided not only support user understanding of model behavior but also offer actionable insights to help users navigate the application and address real-world challenges.
We further validated our classification by examining the AI expertise of recruited participants and the relevance of the AI task to their professional work. As a rule of thumb, studies involving domain experts were more likely to pursue the second objective, since specialists could more easily evaluate the utility of XAI systems in real-world contexts. Conversely, crowdsourced studies were more likely to focus on the first objective, as the tasks were typically less grounded in the specific needs and contexts of real-world users. Of these, we flagged studies that employed attention checks or screening criteria to recruit participants with relevant domain knowledge.
Step 2 - Identifying Design Strategies: During open coding, we first collected and organized prior XAI systems based on different types of explanations (Appendix A). Next, we scrutinized the presentation of explanations, focusing on design components such as visual elements (e.g., lines, shapes, and graphics) and linguistic features (e.g., languages and tones). We inductively derived an initial set of codes characterizing recurring design strategies from this process.
Step 3 - Associating Design Strategies with Rhetorical Appeals: Finally, we conducted axial coding to organize recurring design patterns by aligning them with the three rhetorical appeals: logical reasoning, credibility, and emotional resonance. We analyzed reported quantitative and qualitative user feedback to assess how design strategies influenced user perceptions. This step connected design intent to user impact by examining how specific strategies contributed to rhetorical persuasiveness.
For example, for quantitative measurements, “simulatability” (i.e., participants predicting AI behavior) and “perceived usefulness” assess the effects of logical appeals, as they reflect improvements in user understanding of AI reasoning and AI utility. For credibility appeals, “self-confidence” and “perceived trust” indicate how trustworthy the AI system appears. Similarly, for emotional appeals, “affective reaction scores” and “user engagement metrics” quantify the extent to which explanations resonate affectively with users.
Additionally, we contextualized quantitative measures with reported qualitative feedback. We present the reviewed articles and their corresponding coding schemes in Table LABEL:tab:coding_scheme. For each paper, we specified the primary XAI goal, identified design strategies, and mapped both quantitative and qualitative feedback to the appropriate rhetorical appeals. Each code assigned to an excerpt from the reviewed papers indicates the design intent it represents and the rhetorical appeal it conveys.
However, we acknowledge that our meta-analysis is selective rather than comprehensive. For example, in some cases, design strategies identified from papers primarily targeting the first XAI goal could also be applicable to the second goal. Moreover, not all user studies in prior work captured responses across all three rhetorical dimensions. We further elaborate these limitations in Section 6.3.
5. Results
Drawing on our coding process (Section 4), we report synthesized design strategies from prior XAI research in this section. We categorize strategies based on three rhetorical appeals: Logical Reasoning (Section 5.1), Credibility (Section 5.2), and Emotional Resonance (Section 5.3). Within each appeal, we describe how different design strategies support the two XAI objectives: explaining how AI works or why AI merits use (see Table 1). We assess each strategy’s strengths and limitations based on our axial coding process, linking reported user experiences to their intended rhetorical appeals.
5.1. Logical Reasoning
A key focus of XAI is explaining the logical reasoning behind automated decisions to foster appropriate trust in AI predictions or align user mental models with overall AI behaviors (Wang et al., 2019; Miller, 2019; El-Assady and Moruzzi, 2022). Here, we identify three design strategies from prior XAI research: Isolated Presentation, Adaptive and Mechanistic Presentation, and Graphical Synergy. Explanations can also communicate why AI merits use by positioning the AI as an intelligent, efficient, and rational collaborator. Toward this end, we identify two design strategies that directly enhance the usability and adoption of an AI system: Explanations as Recommenders and Explanations as Retrievers.
5.1.1. Isolated Presentation
Different types of AI explanations (see Appendix A) inherently embody distinct forms of human logical reasoning. Wang et al. (2019) conceptually mapped how various machine learning models reason and generate explanations to corresponding forms of human decision-making. For example, they argued that Bayesian probability represents inductive reasoning, a bottom-up approach that uses prior intuitions to draw broader conclusions. Case-based models encapsulate analogical reasoning, which solves new problems by synthesizing solutions from previously solved similar issues. Rule-based models illustrate deductive reasoning, a top-down approach that begins with general principles and operationalizes them to specific cases.
Explanations can be compelling logical arguments when their embodied reasoning aligns with user mental models. When explaining how AI works, the presented explanation reinforces trust and interpretability if users perceive the mapping from inputs to outputs as logically coherent or aligned with their expectations. In summary, different types of AI explanations operationalize different modes of human reasoning, and their persuasiveness depends on how well embodied logic align with user expectations.
A dominant design strategy employed in prior XAI studies is isolated presentation, where different types of explanations are presented individually for A/B testing. Specifically, these studies examined the extent to which distinct logical structures embodied in different explanations (e.g., features, examples, rules, or concepts) helped to improve human-AI complementary performance (Bansal et al., 2021; Gajos, 2024; Nguyen et al., 2018) or promote appropriate reliance on AI predictions (Schemmer et al., 2023; Yang et al., 2020; Zhang et al., 2020; Liu et al., 2021; Vasconcelos et al., 2023). For example, in comparing feature-based and example-based explanations, Chen et al. (2023b) observed that the latter elicited more critical reflections on potential misalignments between human intuition and machine reasoning. Moreover, van der Waa et al. (2021) demonstrated that rule-based explanation showed a small positive effect on (global) system understanding compared to example-based explanation. Nevertheless, both types of explanations could lead to confirmation bias (Bertrand et al., 2022), where users tended to over-rely on AI predictions erroneously.
Regardless of format or style, all AI explanations are paraphrased distillations of the underlying computational algorithm. Isolated evaluations of explanations may simplify analysis, but can also mislead users by reinforcing prior (erroneous) assumptions or painting an incomplete picture of system behavior (El-Assady and Moruzzi, 2022). These risks arise in part because AI systems and the explainable techniques used to interpret them (described in Section 2.1) are not inherently designed to align with human reasoning (Chen et al., 2023a; Wang et al., 2019). Although neural networks could be loosely modeled on the structure of human brain (Schaeffer et al., 2022), AI systems are typically trained to rely on complex, uninterpretable features that differ fundamentally from the semantic reasoning humans use to understand the world (Fei-Fei et al., 2007; Zhang et al., 2019). As a result, explanations may appear plausibly persuasive to users, but do not necessarily reflect the actual decision-making process of the system or reliably align with human reasoning.
Furthermore, explanations may be misinterpreted when subtle variations in visual elements convey unintended or oversimplified meanings (Lupton and Phillips, 2015; Gibson, 1899). For example, Ye et al. (2023) observed that subtle differences in saliency maps design could significantly influence annotator perceptions in uncertainty-aware medical semantic segmentation. They compared three formats — singular highlights, continuous gradient maps, and confidence contour maps — and found that the first two were less interpretable. In particular, continuous gradients were prone to confusion, as annotators struggled to infer clear thresholds for distinguishing uncertainty levels. Other studies reported similar qualitative feedback, where participants noted that continuous gradient maps of local feature importance only helped them recognize relevant features but made it difficult to quantify their relative importance (Alqaraawi et al., 2020). These failures were also observed in textual contexts (Schuff et al., 2022).
Based on extensive prior research examining the user effect of various explanations, presenting isolated explanations may be most suitable in early-stage design and evaluation settings, where the goal is to investigate how different explanation formats align with the needs (Liao et al., 2020) and preference of a target stakeholder group (Dhanorkar et al., 2021; Panigutti et al., 2023; Tonekaboni et al., 2019). This approach is also effective for formulating user mental models in controlled experiments, such as identifying the error boundaries of human-AI decision-making (Bansal et al., 2019a, b). Notably, isolated explanations provide a clear basis for comparison before progressing to more advance designs, such as adaptive and mechanistic presentation, or graphical synergy, which reveal AI behaviors in greater complexity (as described in the following sections).
5.1.2. Adaptive and Mechanistic Presentation
Since AI systems do not inherently follow human reasoning, presenting isolated explanations risks reinforcing misconceptions by oversimplifying AI behavior. To mitigate this, researchers have adopted two design strategies. Adaptive presentation dynamically adjusts explanation format based on evolving user objectives or deployment contexts. Mechanistic presentation combines multiple explanation types into an integrated UI dashboard. We discuss these two strategies in tandem because both extend isolated presentations through adaptive adjustments or mechanical aggregations. While the mode of delivery differs, the underlying logical appeals of these explanations remain unchanged. This contrasts with synergistic design strategies introduced later (Section 5.1.3), which fundamentally reconfigure explanatory logic.
Adaptive explanations are primarily delivered in two ways: (1) through user requests, such as in a conversational setting, or (2) automatically triggered through internal system indicators. For user-initiated requests, Miller (2019) showcased an idealized explanation dialogue in which an XAI agent provides different types of explanations (e.g., features, rules, examples) in response to diverse user inquiries on how the AI classifies an input image as a spider (versus a beetle or octopus). Shen et al. (2023) operationalized this adaptive dialogue in human-AI scientific writing contexts. They constructed a chatbot to provide heterogeneous AI explanations across different stages of human-AI communication.
For system-triggered deliveries, explanations are automatically adjusted in response to internal (invisible to end-users) indicators. For example, some XAI systems present explanations only when AI confidence is low (Bansal et al., 2021) or dynamically adjust explanations based on real-time accuracy updates or error boundaries (Bansal et al., 2019b, a; Wang and Yin, 2023). Others adapt explanations based on human learning progress (Wambsganss et al., 2020) or show explanations only when the AI’s predicted correctness is estimated to exceed that of the user (Ma et al., 2023).
Compared to adaptive presentations, mechanistic presentations of explanations are generally more straightforward to implement since no real-time adaptations are required. For example, Wang et al. (2019) integrated local feature importance, global feature contributions, and counterfactual rules into an AI-assisted medical diagnosis tool, enabling clinicians to consult multiple explanation types simultaneously when evaluating patient cases. Similarly, Bove et al. (2022) developed an explainable auto-insurance interface that helped vehicle owners understand how their driving behaviors influenced insurance premiums. Likewise, Cau and Spano (2025) created an explainable hiring dashboard that enabled recruiters to accept or reject applications based on applicant attributes presented through multiple concurrent types of explanations.
Overall, both adaptive and mechanistic presentations offer unique benefits to improve user understanding and support human-AI teaming. Adaptive presentations enhance relevance by tailoring explanations to user goals, expertise, or decision context, reducing cognitive overload and aligning responses with moment-to-moment needs. Thus, adaptive explanations are particularly valuable in dynamic or iterative workflows, where user needs may evolve (Suresh et al., 2021). Mechanistic presentations offer breadth and transparency by juxtaposing multiple explanation types in a unified interface, allowing users to cross-reference local and global reasoning or compare factual and counterfactual views. As a result, they are more effective in diagnostic or investigative tasks, where users benefit from cross-checking different explanatory lenses to reach more robust conclusions (Miller, 2023a). Compared to isolated presentations, interactivity is a crucial differentiating factor for adaptive and mechanistic presentations, allowing logical reasoning processes to unfold dynamically in response to user actions or system feedback.
Despite their benefits, Gaole et al. (2025) found that both adaptive and mechanistic presentation formats could lead to a clear over-reliance on the AI system. The authors attributed this over-reliance to an illusion of explanatory depth where the presence of multiple explanations, regardless of actual fidelity, reduced critical evaluation of system behavior. This echoes a classic rhetorical strategy in communication, where shifting the burden of proof makes arguments appear more persuasive (Hahn and Oaksford, 2007).
5.1.3. Graphical Synergy
Regardless of whether explanations are presented in isolated, adaptive, or mechanistic formats, different explanation types often coexist unilaterally without being cohesively integrated. As Wang et al. (2019) noted, this fragmentation arises because the underlying explanatory algorithms rely on fundamentally different logical bases. Nevertheless, emerging work has shown promise in organically merging diverse explanation types into a visually coherent representation, an approach we term graphical synergy. Due to its visual nature, graphical synergy can be effective for pattern recognition tasks, where synthesizing complementary explanation types helps users grasp subtle relationships that no single explanation could convey.
For example, Evans et al. (2022) synergistically combined normative example-based explanations with counterfactual explanations, forming a cohesively contrastive causal chain. Specifically, they presented prototypical examples alongside counterfactual instances of cell nuclei to help clinicians differentiate between positive and negative Ki67 scores used to estimate tumor proliferation. Compared with traditional saliency maps, prototypical examples, and concept attributions, clinicians reported that the synergistic presentation “help[ed] them better understand what the algorithm is looking for… it is self-explanatory that staining was the most important factor distinguishing positive nuclei from those marked negative, with some identifying other important factors such as shape [and] size of nuclei.” Clinicians’ subjective ratings reinforced this feedback, where the synergistic explanation format received the highest median scores for clarity, relevance, trustworthiness, and practical utility.
While graphical synergy can improve explanation coherence, integrating multiple explanation types can raise concerns about fidelity to the underlying AI model. To synthesize transitional examples between prototypical positive and negative Ki67 nuclei instances, Evans et al. (2022) combined data attribution methods with image morphing techniques (Gulshad and Smeulders, 2021). Because these synthetic interpolations are not grounded in training data, they raise faithfulness concerns that highlight the need for new XAI techniques that seamlessly integrate explanation types while preserving fidelity. Developing such techniques would support richer graphical synergies, such as combining example-based explanations with saliency highlights (Kenny and Keane, 2021) or integrating salience maps with human-interpretable concepts (Yang et al., 2024).
5.1.4. Explanations as Recommenders
Traditionally, XAI work focused on explaining how systems work by helping users to assess, replicate, or predict AI decisions. Emerging work now shifts toward explaining why AI systems merit use, particularly by supporting users in solving real-world problems beyond simple classification. In these contexts, explanations act as recommenders, helping users understand what actions to take and why those actions are beneficial.
For example, Wambsganss et al. (2020) developed an adaptive AI-assisted learning system to help students write persuasive arguments. To facilitate revisions, the system visualized argumentative structure, evaluated writing quality, and provided natural language feedback alongside highlighted areas for improvement. Students expressed a significantly greater preference for using the new system than traditional reference-based writing tools. Nilforoshan and Wu (2018) implemented a similar application using feature-based explanations to improve reviewer responses on social media.
Besides writing, algorithmic recourse represents another practical use case of explanations. For instance, Kim et al. (2024) developed an XAI interface that combined counterfactual examples with recourse plans to help dysarthria patients correct mispronunciations. Similarly, Wang et al. (2023b) introduced GamCoach, which provided loan applicants with counterfactual explanations recommending financial changes to improve loan approval odds.
Findings from Wang et al. (2023b) highlight a key point that aligns with our framework (outlined in Section 3.3): effective use of an AI system does not necessarily require a deep understanding of system behaviors. For example, in their user study, participants rated “ease of understanding” (4.90) and “help with understanding the model” (5.17) lower than their “intention to reuse the system for future loan applications” (5.34) and “the system’s usefulness in identifying ways to improve their loan applications” (5.61), based on a 7-point Likert scale. Although observed differences were not statistically validated, this finding suggests that a deep understanding of model behaviors is not a prerequisite for effectively using the AI for algorithmic recourse.
To summarize, explanations as recommenders are particularly beneficial in scenarios where users seek actionable guidance. By positioning AI as a collaborative advisor, this strategy reinforces our position on XAI’s rhetorical role: explanations should reveal how AI works and why an AI merits use for practical utility.
5.1.5. Explanations as Retrievers
Explanations can also function as retrievers to support dataset exploration and analogical reasoning. Different types of explanations highlight distinct model signals (see Appendix A), including local feature attributions, global patterns, representative prototypes, and human-interpret concepts. Users can leverage these characteristics to retrieve analogous cases, facilitate structured comparisons along shared attributes, and derive actionable insights grounded in prior experience. Building on this idea, researchers have developed interactive case-based retrieval systems that surfaced prior instances based on similar features or concepts through explanations. Applications include clinical diagnosis review (Cai et al., 2019b; Suresh et al., 2022), auditing dataset annotation quality (Kenny et al., 2021), identifying model vulnerabilities (Sun et al., 2023a), and refining chatbot behaviors (Miranda et al., 2021).
Consider Cai et al. (2019b)’s case-based retrieval tool for medical diagnosis. Rather than evaluating AI outputs directly, pathologists used the tool to refine search queries that retrieved similar scans with confirmed diagnoses, supporting their interpretation of the current image. Pathologists improved queries by either cropping salient regions from normative scans or adjusting the feature importance of known medical concepts. Here, the primary XAI objective is to improve diagnostic accuracy for pathologists (i.e., why AI merits use) rather than understanding and verifying AI predictions (i.e., how AI works).
5.2. Credibility
Explanations can influence users’ perceived trust in AI systems when designed to project credibility. For the first goal of understanding how AI works, we identify three credibility-focused design strategies: Authority Framing, Cultural Relatedness, and Numerical Association. To communicate why an AI system is trustworthy and worth adopting, we identify two additional strategies: Technical Assurance and Social Endorsement.
5.2.1. Authority Framing
HCI researchers have examined how different communication styles affect user acceptance of machine-generated advice. These styles include direct versus implicit recommendations (Rau et al., 2009) and languages that signal authority or expertise (Metzger et al., 2024; Saunderson and Nejat, 2021). Naturally, designers have integrated authoritative language and tone into explanations to enhance their perceived credibility (Cambria et al., 2023b; Okoso et al., 2025). We refer to this design strategy as Authority Framing.
For example, an AI-recommended insulin dose can be phrased persuasively to make patients more likely to trust and act on the recommendation. In a detached or impersonal tone, the technical explanation might read:
“Because your planned alcohol intake exceeds one unit, the system recommends a lower dose of insulin. A normal dose would have been appropriate if the intake were one unit or less.”
(Example adapted from van der Waa et al. (2021))
Alternatively, the same explanation can be delivered in a more supportive tone, echoing a nursing communication style that balances empathy with expertise.
“Planning to enjoy more than one drink? The system suggests making a small adjustment to your insulin dose to keep you feeling your best. If you only have one unit or less, your usual dose is just right. We’re here to support you in making healthier choices!”
Another example is an AI tour guide by Díaz-Rodríguez and Pisoni (2020), which combined expressive language with domain expertise to explain the historical context of paintings, sculptures, and other artworks. Visitors responded more positively when explanations conveyed both factual accuracy and emotional resonance. Thus, authority framing might be most effective for personal contexts, where users interpret and act on system guidance that directly affects their behaviors. In these cases, combining expertise with empathy not only increases the perceived credibility of the recommendations but also makes users feel emotionally supported in making sensitive or value-laden decisions (Okolo et al., 2024; Komatsu et al., 2020).
5.2.2. Cultural Relatedness
Beyond linguistic styles, XAI researchers have also integrated visual elements (e.g., color and shape) into explanations to project a credible tone, voice, or personality. Users are more likely to understand and trust explanations when presented visual elements resonate with their beliefs, values, or cultural identities. We refer to this design strategy as Cultural Relatedness.
For example, Okolo et al. (2024) studied the deployment of a neonatal jaundice [^3] Neonatal jaundice is a prevalent health condition in newborns, marked by yellowing of the skin and eyes. detection application in rural Indian clinics. During a co-design workshop, community health workers raised concerns about the default color scheme used in LIME (Ribeiro et al., 2016) and SHAP (Ghorbani and Zou, 2019) visualizations. In these explanatory visualizations, red simply denoted a classification outcome. However, health workers found the red color schemes confusing because red typically signifies severity and danger in medical contexts. To address this concern, workers applied traditional Hindu colors and styles to replace default color schemes in LIME and SHAP.
To avoid misinterpretations associated with colors, Okolo et al. (2024) experimented with other visual techniques to convey feature importance, such as textual descriptions and different shapes (e.g., squares and circles). Clinicians found that replacing color with text improved clarity, but using shapes caused new confusion, as some interpreted quantity (e.g., number of circles) as a proxy for condition severity. These findings emphasize the need for culturally informed visual designs to avoid unintended miscommunications.
5.2.3. Numerical Association
Another common technique to project credibility is numerical association, where explanations embed numbers, scores, or other metrics to reinforce perceived objectivity and authority. Numbers often carry a rhetorical power: they signal precision, rigor, and algorithmic sophistication. Naturally, numerical association is effective in quantitative contexts (e.g., performance reviews, financial scoring, or risk assessments), where users may benefit from the perceived objectivity of numbers to legitimize, clarify, and compare AI recommendations. On the other hand, users may also mistakenly equate numerical information with higher system competence, even if they do not fully understand its statistical meaning (Li et al., 2023).
For example, Ehsan et al. (2024a) implemented three different styles of AI explanations to evaluate how users perceive an AI agent’s intelligence. Regardless of participants’ backgrounds in AI knowledge, users consistently rated the agent with numerical explanations as more intelligent and reliable, even when they lacked a clear understanding of the presented metrics. The authors attributed this unwarranted faith to preconceived associations between numerical information and algorithmic thinking processes, reinforcing a perception of objectivity without true interpretability.
These examples highlight a key challenge: managing the tradeoff between credibility (ethos) and logical clarity (logos). While design strategies like cultural relatedness and numerical association can increase users’ perceived trust in AI systems, they do not always support a better understanding of how the system works. In some cases, as observed in Okolo et al. (2024) and Ehsan et al. (2024a), credibility-focused strategies may even obscure how AI systems work by introducing misinterpretations or inducing false confidence. See Section 6.3.1 for further discussion on responsibly signaling authority.
5.2.4. Technical Assurance
In practice, AI developers often employ technical criteria to validate predictive accuracy and fairness. Examples of technical standards include mitigating model bias (Gupta et al., 2023) and ensuring robustness against adversarial inputs (Gupta et al., 2024). Incorporating these criteria into explanations can foster perceived user trust by clarifying the AI system’s legitimacy and justifying why it merits use. For instance, Mitchell et al. (2019)’s model cards and Ribeiro et al. (2020)’s model behavior checklists exemplify how technical reliability can be assessed and communicated to users. Their design strategies convey an implicit message of assurance: We have made every effort to ensure this AI model is reliable so that you can use it safely and confidently.
While primarily for developers, tools like model cards and behavior checklists also serve a broader purpose. They signal that the model aligns with rigorous technical standards, which can subtly influence perceived user trust and adoption in practice. Like car manuals, which are rarely used but still inspire confidence, the real value of these technical artifacts lies in conveying assurance and reliability. Moreover, these artifacts also help brand a company’s identity around commitments to quality and responsible innovation.
To support complementary human-AI teaming, designers have incorporated technical criteria alongside adaptive explanations (introduced in Section 5.1). For example, Zhang et al. (2020) incorporated AI confidence scores alongside feature-based explanations to examine whether displaying confidence levels would increase users’ likelihood of accepting AI predictions. Similarly, Bansal et al. (2021) adapted the presentation of explanations based on the system’s confidence, showing explanations only when the model’s predicted confidence was high. Both studies found that participants heavily relied on these technical signals to guide their acceptance of AI predictions.
However, when exposed to technical indicators such as confidence scores, users may blindly follow AI predictions without critically assessing explanations. Despite this concern, researchers have yet to identify the most effective strategy to mitigate over-reliance. Some researchers have explored adding cognitive forcing functions to encourage more critical engagements with AI explanations (Gajos and Mamykina, 2022). Other researchers attribute the functional utility of explanations to domain experts, suggesting that their effectiveness depends on the user’s prior knowledge (Bansal et al., 2021; Wang and Yin, 2021). These works highlight a key trade-off between AI explanations’ functional utility and persuasive credibility. If users ultimately learn to perform the task independently, what added value does the AI system continue to provide?
We argue that engineers and technical developers should be responsible for ensuring that technical criteria support the responsible use of AI. Given the limited expertise of lay users, it is unrealistic to expect them to critically evaluate the reliability of technical signals. Once technical signals can faithfully convey the AI’s reliability, they should be carefully integrated into user-facing designs to promote assurance and confident use.
5.2.5. Social Endorsement
Beyond incorporating technical assurances, researchers have also employed social endorsements to strengthen users’ perceived trust in AI systems. In organizational settings, social affirmations may exert a greater influence than the explanations themselves on whether users accept or reject AI decisions.
For example, Ehsan et al. (2021) developed an explainable AI interface for sales recommendations. Alongside feature-based explanations, the system provided a historical trajectory that showed how past sales teammates responded to the AI recommendation and the rationales behind their choices. Participants reported that these design elements enhanced their sense of social validation, transitive trust, and the temporal relevance of AI recommendations within the organizational setting.
While Ehsan et al. (2021) framed their design as promoting sociotechnical transparency, we interpret their approach as an example of social endorsement. Similar to how product reviews influence consumer behavior (Talib and Saat, 2017), their design leverages peer adoption to foster trust. When new users see that trusted colleagues have adopted and endorsed the system, they are more inclined to perceive the AI as credible and worth using. Social endorsements can also signal alignment with institutional norms, allowing trust to propagate socially through transitive peer relationships and collaborative culture.
5.3. Emotional Resonance
Emotion has historically played a crucial role in design (Ho and Siu, 2012; Norman, 2007), serving essential communicative and experiential functions. In XAI, researchers have incorporated affective elements into explanations to make them more engaging. We identify three design strategies to support explaining how AI works: Explanation Affect, Emotional Cues, and Protective Framing. Additionally, we identify two strategies to explain why AI applications are emotionally compelling to use: Expectation Management and Expressive Interaction.
5.3.1. Explanation Affect
Different types of AI explanations (see Appendix A) can inherently evoke varying emotional responses depending on their relevance to the task, a phenomenon known as explanation affect. For example, Bernardo and Seva (2023) implemented example-based, feature-based, and rule-based explanations for animal identification. Participants judged the emotional appeal of each explanation type, highlighting differences in how each form engaged users. The study reported that example-based explanations elicited stronger feelings of surprise, trust, and perceived usefulness. In contrast, feature-based and rule-based explanations led to reduced emotional engagement.
However, the authors cautioned that these findings might not generalize to other tasks. In their study, the red color used in feature-based explanations to mark the Tarsier’s eyes and ears might have unintentionally evoked negative emotions such as fear or discomfort, an effect absent in the presented example-based explanations. Additionally, the author emphasized that users’ prior emotional states and attitudes toward AI could strongly influence how different explanation formats affect decision-making and trust.
Thus, in practice, explanation affect emphasizes matching the emotional fit between the explanation form and the underlying human concern. In public-facing applications, such as wildlife conservation or sustainability, example-based explanations may be particularly effective, as they evoke empathy and foster personal connections by referencing real-world past cases. In contrast, in more high-stakes settings, such as medical treatment (van der Waa et al., 2021), recidivism screening and financial decision-making (Wang et al., 2022), rule-based explanations are preferable due to their capacity to convey structure, fairness, and procedural consistency.
5.3.2. Emotional Cues
Unlike explanation affect, where different types of explanations inherently evoke diverse emotions, some XAI researchers have directly integrated emotional cues into explanations to enhance user engagement. Examples of emotional cues include emojis, color, and tone of language.
For instance, Freire et al. (2023) incorporated emojis into AI explanations to increase user engagement. Bernardo and Seva (2023) compared robotic versus human-like language styles and found that robotic phrasing heightened anxiety and reduced perceived usefulness, while human-like language eased fear and improved perceived utility. In some cases, however, emotional responses are triggered unintentionally. For example, Okolo et al. (2024) found that the red color used in LIME and SHAP visualizations evoked negative emotions in a medical context, where red is commonly associated with danger, even though this was not the intended message of the explanation.
Additionally, researchers have incorporated emotional cues into interactive AI explanations. For example, in Wang et al. (2023b)’s GamCoach (introduced in Section 5.1.4), the design leveraged different colors, emojis, and visual cues to recommend interactions for algorithmic recourse. Participants found the tool relatively easy and enjoyable to use.
As emotional cues draw on visual, linguistic, or stylistic elements to elicit affective engagement, these characteristics can be particularly effective in educational AI tools. When integrated with adaptive explanations, as shown in the augmentation-based writing systems developed by Wambsganss et al. (2020), emotional cues can enhance memory retention and spark curiosity, thereby enriching students’ learning experiences.
5.3.3. Protective Framing
Emotional cues are typically designed to elicit positive reactions toward AI explanations, making them more engaging, approachable, or enjoyable. In contrast, protective framing [^4] We adopt this term from Fokkinga and Desmet (2012), who explored how people use negative emotions in everyday life to achieve goals relies on negative emotions like skepticism or caution to encourage more critical reflection of AI explanations. Our narrative literature review did not identify design examples that explicitly employed this strategy. Instead, many studies introduced algorithmic aversions or AI system failures as independent variables to examine user acceptance of AI decisions and engagement with explanations. Thus, we propose protective framing as a possible direction for future XAI design.
For example, Ebermann et al. (2023) and Ben David et al. (2021) found that negative feelings such as disappointment, annoyance, or frustration often arise when users encounter unexpected AI outcomes or failures. These adverse emotions can heighten users’ attention to explanations. Notably, Ben David et al. (2021) observed that following an AI failure, participants were significantly less likely to accept subsequent AI advice when explanations were shown, compared to when no explanations were shown. This suggests that explanations can amplify critical scrutiny under negative affect.
Building on this observation, we could imagine a design scenario illustrating how protective framing could facilitate more critical user engagement. When the AI prediction contrasts with a user’s expectations, a system alert might state:
“The AI has predicted a different outcome, but it’s important to remember that both humans and AI can sometimes make mistakes. We encourage you to review the detailed explanation behind the AI’s decision…[show AI explanations]”
Even when the AI prediction aligns with the user’s decision, the system can still deliver a cautionary alert, reminding users of the potential for AI to make errors and encouraging them to review the explanation critically. This cautionary disclaimer has been shown to act as a cognitive forcing function, effectively slowing down the decision-making process and prompting users to process explanations more carefully (Gajos and Mamykina, 2022; Buçinca et al., 2021).
For practitioners, protective framing helps directly safeguard against overreliance on AI systems. Specifically, failures could serve as a protective frame, prompting users to explore AI explanations proactively before finalizing their decisions. This idea reflects an alternate interpretation of the cost-benefit model described by Vasconcelos et al. (2023). While their work focused on reducing the cost or increasing the benefits of verifying AI predictions (through explanations), triggering negative emotions emphasizes the consequences of neglecting verification, effectively increasing the perceived benefit of careful review.
5.3.4. Expectation Management
In real-world AI applications, explanations are commonly used to manage user expectations. Before interaction, explanations can help justify why the AI system is worth using by outlining its strengths and limitations under intended use cases. After interaction, explanations clarify how the AI works by outlining the factors influencing its recommendations. This mirrors how people use explanations to set and adjust expectations, maintain shared understanding, and resolve ambiguity in everyday communication.
For example, Kocielnik et al. (2019) found that presenting AI explanations upfront could lower user expectations of AI performance, thereby reducing dissatisfaction during AI breakdowns. However, this emotional remedy may not be effective when explanations are not perceived as practically helpful. For example, Ebermann et al. (2023) found that participant mood scores significantly declined when explanations justified alternative decisions rather than the AI’s predictions, creating cognitive dissonance and undermining user satisfaction. Interestingly, despite the observed negative emotional impact, participant evaluations of system usefulness (i.e., self-rating on the question “to what extent was my [AI] support helpful”) actually improved. This divergence suggests that emotional discomfort does not always indicate poor system design. Like how students may resist challenging coursework yet still benefit from it, friction in human-AI interaction may play a productive role by encouraging more critical system evaluations.
For practitioners, expectation management can calibrate user expectations, preempt dissatisfaction, and maintain trust when introducing a novel AI system under early development (i.e., when system performance may be unstable). Additionally, expectation management can help mitigate algorithmic aversion (Hou and Jung, 2021; Cheng and Chouldechova, 2023) by clarifying an AI system’s strengths, limitations, and appropriate boundaries of use, making the system appear more transparent and aligned with users’ concerns about overreach or misuse.
5.3.5. Expressive Interaction
Expectation management primarily treats explanations as affective band-aids to remediate AI breakdowns. While this reactive use of explanations can be effective, designing with emotional elements also calls for a more proactive and holistic perspective. Beyond remediation, AI explanations can enable expressive interaction, a design approach that sustains emotional engagement throughout the user’s interactive journey with the AI system.
For example, Alvarado et al. (2022) developed a humanoid controller to interact with a movie recommendation algorithm. Users can accept or reject AI-suggested movies through various physical controls, such as pressing, holding, or shaking the controller. Notably, a long press at the controller’s center triggers the system to explain why a particular movie was recommended based on the user’s viewing history or personal preferences. This gesture provides a subtle way to address user curiosity about why a particular recommendation was made. More importantly, AI explanations are woven into a larger embodied interaction, functioning as a natural part of the experience rather than as isolated add-ons.
Expressive interaction treats emotion not as a remedy, but as a generative force that actively shapes how users engage with AI systems. Rather than smoothing over frictions, this design strategy aims to seamlessly integrate emotional responses as a natural part of human-AI interplay. Through repeated interactions, users develop personal rhythms with the system: deciding when to seek explanations, how much to rely on them, and what emotions to attach to different types of feedback. In this way, explanations become part of an evolving emotional narrative, rather than static tools for understanding model behavior. As a result, expressive interaction naturally supports long-term engagement applications by enabling personalized explanation experiences that adapt to users’ evolving needs, emotions, and goals.
6. Discussion
By synthesizing design strategies employed in prior XAI systems, we have identified a pluralistic design landscape, reflecting how researchers have explained how AI works and why AI merits use. Since the reviewed literature captured only a subset of existing design strategies, our synthesis should be viewed as a starting point rather than a comprehensive catalog. As XAI evolves, new strategies will likely emerge for different rhetorical appeals. In this section, we discuss the contributions of these diverse strategies to existing XAI research (Section 6.1 and 6.2), the caveats of identified design strategies, and limitations of our review process (Section 6.3).
6.1. Centering XAI Around User Intent and Task Relevance
As described in our rhetorical framework (Section 3.3), we extend the traditional goals of XAI beyond merely clarifying model behavior to include another dimension: explaining why AI merits use. Note that this goal is better understood as a design objective rather than a technical optimization. When explaining why AI merits use, it is important to consider when AI is appropriate, when it is not, and the degree of suitability for a task. Factors such as model reliability (Petch et al., 2022), potential consequences of errors (Ehsan et al., 2022a), and alignment with user goals help users make informed adoption decisions. As highlighted in Section 5, prior researchers have already employed various design strategies to support this objective. Some used explanations as recommenders or retrievers, offering actionable insights for real-world problem solving. Others applied credible or emotional framing to make AI applications more compelling and persuasive.
A natural direction for future XAI research is to incorporate user intent, task relevance, and model reliability more explicitly as design criteria when selecting and evaluating explanations. In particular, researchers might consider: 1) whether a user’s strong motivation to adopt AI for a given task creates a meaningful need for explanations, 2) to what extent that need varies by task contexts, and 3) how reliably the AI can be used in conjunction with explanations that promote safe and informed decisions. Connecting user intent, AI robustness, and explanatory relevance in this way can enhance both the interpretive value and empirical rigor of XAI research.
For example, when examining the effects of different AI explanations on sentiment classification, recidivism prediction, or other tasks, the first set of questions could start from: Who are the targeted users of these AI tasks, and why would they want AI to assist them? Answers to these questions could provide specific use cases to further investigate users’ explanatory needs: Why would they care about understanding individual predictions or model global behavior in these tasks? What AI explanations and design strategies could better achieve their goals for using AI? Even when traditional AI explanations fall short, answering these questions can reveal new user needs and prompt designers to rethink the problem space, leading to more relevant solutions (Zimmerman et al., 2022). This shift in perspective opens opportunities to repurpose XAI techniques to support emerging AI applications beyond their original intent (Liu et al., 2024).
6.2. Perceiving Explanation Design as a Reflective Rhetorical Practice
In HCI, there is growing interest in promoting stakeholder participation and integrating cross-disciplinary knowledge when developing XAI systems (Malizia and Paternò, 2023; Mohseni et al., 2021; Dhanorkar et al., 2021). Within these collaborative settings, designers can play a critical role in addressing diverse stakeholder needs by drawing on the strategies synthesized in Section 5. Specifically, designers could balance different rhetorical appeals to accommodate the diversity and flexibility of human needs. For example, in some use cases, logical comprehension of the AI system and its practical utility take precedence, with credibility and emotional factors serving a secondary role. This is often the case when designing AI-powered analytical tools for technical users, such as developers or data scientists. In contrast, for consumer-facing AI applications, credibility and emotional resonance may be more influential in enhancing user engagement and trust.
Additionally, different design strategies can help mitigate one another’s limitations. For example, Gaole et al. (2025) observed that users tended to overrely on the AI system when presented with conversational or mechanistic explanations, which created an illusion of explanatory depth. To address this, designers may incorporate emotional strategies such as protective framing or expectation management to encourage more critical user engagement.
However, integrating and balancing multiple design strategies to meet diverse stakeholder goals remains a complex, evolving challenge requiring future work. For example, an XAI design solution that worked well at one point, or for a particular stakeholder group in a specific domain, may become inadequate as the application context evolves, necessitating new design responses (Friedman, 2008; Zimmerman et al., 2022). Buchanan (1992) argued that design is a form of inquiry, where the actual problems emerge through continuous user engagement, reflection, and iterative assessment. This perspective is particularly relevant for designing AI explanations, which increasingly involves collaborative and sociotechnical considerations (discussed in Sections 2.3 and 2.4). Thus, addressing this challenge may require designers to use an iterative, exploratory design process (e.g., research through design (Zimmerman and Forlizzi, 2014)) combined with creative strategies to navigate potentially competing and shifting needs.
6.3. Rhetorical Design Risks and Limitations of Our Review Process
6.3.1. Misuses of Design Strategies
While rhetorical strategies can make AI explanations more compelling, they inadvertently carry ethical risks. If misapplied without guardrails, these strategies can blur the line between persuasion and manipulation, raising concerns about transparency, fairness, and user autonomy. Some risks were already discussed when we introduced each rhetorical strategy. Therefore, in this section, we briefly outline some broad patterns of potential misuse across the three rhetorical dimensions (logos, ethos, and pathos). This is not intended as a comprehensive typology, but as a starting point for caution and reflection. As rhetorical XAI becomes more widely adopted, we recommend a deeper analysis of how to balance persuasive design with responsible practice as future work.
Logos One risk pattern in logos is the use of selective logical appeals to create explanatory fallacies, where explanations appear logically coherent but are actually incomplete or misleading. These fallacies may arise from intentional disinformation or unintentional misinformation conveyed in the explanations. For example, explanations as recommenders (5.1.4) may only present advantageous factors that support adoption while omitting adverse side effects. Consider a clinical decision-support system that might explain its recommendation for a new medication by highlighting improved recovery rates, while failing to mention common side effects such as nausea or fatigue. As a result, explanatory fallacies may conceal relevant model reasoning and downplay potential risks, undermining interpretive fidelity (Morrison et al., 2023). Additionally, the decision-support system itself could accidentally omit vital information or hallucinate false information due to limitations in its training data. Thus, it is important for clinicians to actively involve in this process to convey these explanations that is accurate and understandable (Amann et al., 2020).
Ethos Credibility appeals may be deceptive when the source of trust is unverified, unearned or a previously trusted authority loses trust. Strategies like authority framing (5.2.1) or social endorsement (5.2.5) may leverage institutional branding, peer behavior, or delayed fact-checking to signal legitimacy, regardless of actual system quality. For instance, a university admissions algorithm might present its decisions with a badge like “Certified by Admissions Experts,” even if such certification has not undergone verification. By applying the language and visual cues of credibility, unethical AI systems may foster misplaced trust by masking the absence of verified authority.
Pathos For pathos, emotional cues (5.3.2) risk manipulating user behavior when they reinforce decisions that may not align with users’ goals or best interests. For instance, a personal finance assistant may use cheerful animations or green thumbs-up icons to reward users for complying with its savings advice. While such cues foster positive affect, they may also encourage users to accept AI recommendations uncritically, reinforcing confirmation bias (Mercier and Sperber, 2011; El-Assady and Moruzzi, 2022). This illustrates how affective nudges can substitute emotional reassurance for critical thinking, deceptively fostering unwarranted trust without genuine understanding.
These risks of manipulation recall the analogy of designers as rhetoricians (introduced in Section 3.3), a role increasingly associated with negative connotations in marketing (Miles and Nilsson, 2018), adversarial propaganda campaigns (Farrell and Schneier, 2018), and foreign influence operations (Starbird et al., 2019). However, our aim is not to promote rhetorical strategies as a tool for manipulation, but to emphasize the ethical responsibilities designers should uphold when employing these strategies to support stakeholder goals. Following Buchanan (Buchanan, 1985a, b), we position rhetoric as a means to support reasoned invention and explore meaningful design alternatives. At the same time, we recognize that any persuasive strategy carries the risk of misuse, as described above. Introducing rhetorical approaches in explanation designs obliges us and future designers to take responsibility for anticipating how such methods could intentionally or inadvertently mislead or deceive. Future research should not only explore how rhetorical design strategies could be balanced with ethical concerns, but also critically examines safeguards, accountability mechanism, and the contexts where these strategies are most appropriately applied. One approach could involve framing them as dark patterns (Gray et al., 2018; Kran et al., 2025) to warn future XAI designers.
In sum, by reconsidering the role of rhetoric in explanation design, we argue that design strategies are not inherently benign or deceptive; rather, their ethical impact largely depends on how they are employed.When grounded in human-centered values and established principles of responsible communication, such as sustainability (Blevis, 2007), dignity (Kim, 2021; Buchanan, 2001b), and social inclusiveness (Costanza-Chock, 2020; Thorpe and Gamman, 2011), rhetorical explanations can support ethical engagement with AI systems and mitigate risks of misuse. This perspective frames rhetorical strategies as communicative tools that can be informed by the wider canon of communication practices, highlighting the responsibility of designers in shaping their applications.
6.3.2. Limitations of Our Review Process
Future work can refine our literature review process as rhetorical design gains broader recognition in the XAI community and new design strategies are proposed and systematically evaluated.
In our review, we categorized papers according to two primary XAI goals: (1) explaining how AI works, and (2) explaining why AI merits use. We observed that many studies targeting the first goal also included measurements relevant to the second, such as perceived AI usefulness and user trust in the AI system. For example, technical assurance was frequently incorporated in XAI studies to measure how users perceived different explanations, even though these studies did not explicitly articulate user intent and the task relevance (e.g., why participants might use AI to complete the task). Nonetheless, our axial coding incorporated both quantitative and qualitative findings from prior studies to understand the influence of different design strategies in promoting the AI’s benefits. Future reviews might consider more stringent inclusion criteria to ensure more precise alignment between study design, user intent, and task relevance.
Additionally, as described in our axial coding process (Section 4), we mapped user study measurements from prior work to the three rhetorical appeals in our framework (Section 3.3). However, this mapping process was inherently conceptual and subjective, incorporating author biases. On the other hand, this mapping also suggests an opportunity to develop a comprehensive set of measurement tools explicitly aligned with the dimensions of logical reasoning, credibility, and emotional resonance. Future research can systematically examine how existing evaluation metrics align with each rhetorical appeal, followed by a scoping review to understand the effects of different design strategies better.
6.3.3. Practical Implementation Gaps
In this paper, we have introduced a novel taxonomy to organize and reflect on rhetorical strategies in XAI, encouraging researchers to consider designing explanations to not only explain how AI works but also justify why AI merits use. For each strategy, we highlighted potential applications and settings where these approaches may be beneficial. However, determining which strateg(ies) to apply in a given scenario remains an open challenge. Triage and selection will likely depend on context-specific factors and we leave this important direction for future work.
7. Conclusion
In this paper, we propose conceptualizing AI explanations as rhetorical artifacts, designed not only to support user understanding of AI systems but also to promote appropriate persuasive adoption as another dimension of responsible AI. Drawing on the concept of rhetorical design, we develop a rhetorical framework that guides our coding process for a narrative literature review. Through this review, we synthesize a range of design strategies used in prior XAI studies to make explanations logically convincing, credibly appealing, and emotionally engaging. By critically examining the strengths and limitations of these strategies, this paper extends the design space of XAI by providing designers with new tools to address the complex needs of diverse stakeholders and the broader sociotechnical challenges involved in building XAI systems.
Acknowledgments
We thank Ahmer Arif, Jacek Gwizdka, James Howison, Yujin Choi, and Yiwei Wu for their valuable comments. This research was supported in part by Good Systems [^5] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsors.
Appendix A Types of AI Explanations
We briefly summarize AI explanations designed in a static form, including example-based, feature-based, counterfactual, rule-based, and concept-based explanations. These explanations serve as the basic design unit for more complex and interactive AI explanations (Bertrand et al., 2023). Note that this is a design taxonomy based on a logical understanding of how AI works, which might not closely align with the taxonomy used by AI developers, who might prefer the classification of XAI based on interpretable techniques (see Section 2.1).
A.1. Example-based Explanations
Example-based explanations generally refer to the XAI methods that use specific instances that are either retrieved from the training dataset or generated as new data points to explain how AI makes a particular decision. As Molnar (2020) and Wang et al. (2019) emphasize, the benefit of these explanations lies in how humans often use examples or analogies to aid in decision-making.
Researchers have explored different kinds of examples to explain how AI works. The most commonly selected examples are those that share similar features to the input. We refer to these examples as nearest-neighbors examples because “nearest neighbors” are both intuitive and widely used as a common technique for researchers to retrieve such examples. The classes of these nearest-neighbors examples can vary to the predicted output class depending on whether nearest neighbors are directly used in the model influence function. For example, if the prediction model is a k-nearest neighbors model, the examples should share the same output class as the input (Papernot and McDaniel, 2018; Su et al., 2023). However, if other prediction models are used and nearest neighbor is only used as a retrieval technique, the output class of these examples might differ (van der Waa et al., 2021; Cai et al., 2019a; Yang et al., 2020; Wang and Yin, 2021; Chen et al., 2023b). Therefore, it’s important to distinguish these two use scenarios to communicate how AI works. Using nearest neighbors as a retriever can only be interpreted as using examples to examine AI predictions. It might be incorrect to claim that these examples directly influence AI predictions.
Cai et al. (2019a) describe another two kinds of examples based on the class difference of nearest-neighbors examples. They are comparative examples, examples similar to the input but with alternative output classes, and normative examples, prototypical examples with the same class to the input. Both yield a different reasoning process for humans to map AI inputs to outputs based on examples. The idea of comparative examples is similar to counterfactual explanations, informing us how an instance could be changed to alter the AI prediction (Molnar, 2020; Verma et al., 2024) (We elaborate this in Section A.3, counterfactual explanations). In contrast, a normative example serves as a benchmark, allowing people to compare the input with a prototypical example to understand how their similarities result in the same output class.
A.2. Feature-based Explanations
Feature-based explanations generally refer to showing feature attribution or influence by indicating which input feature is important to the model global behavior or how it has direct positive or negative influence towards different output classes (Molnar, 2020; Wang et al., 2019; Mohseni et al., 2021). Unlike example-based explanations that are mostly local interpretations (i.e., how a specific input affects the output based on examples), feature-based explanation can communicate either local or global interpretation (i.e., how AI behaves specifically in a particular instance or generally on all features) depending on using which XAI techniques (Molnar, 2020).
Local feature contribution, which is also named as feature salience, highlights input features showing their contribution for different output classes. In other words, it tells us how each feature directly affects a model to make a negative or positive prediction for the test instance (Ribeiro et al., 2016; Molnar, 2020). In contrast, global feature importance or named feature relevance, would rank input features based on their overall importance that influence the model behavior at large (Lundberg and Lee, 2017; Molnar, 2020). In other words, it indicates how the model relies on different features for making future predictions in general. These two types of feature-based explanations work in a complementary fashion to communicate how AI works based on feature attributions.
A.3. Counterfactual Explanations
Counterfactual explanations characterize a hypothetical scenario telling us what could have happened if something changes (Verma et al., 2024). By adapting this scenario to describe the model’s input-output relationship, counterfactual explanations help inform the smallest change to the input values that alter the model prediction (Verma et al., 2024). Molnar (2020) highlights that the presentation of counterfactual explanations are always example-based as they have to be different instances showing how to change the test instance.
Based on the above definition, the differences between counterfactual explanations and several types of example-based explanations described earlier could become much clearer. First, we can infer that comparative examples are always counterfactual but counterfactual explanations do not need to be actual examples existed in the training data. They can be hypothetical examples by changing feature values. More importantly, counterfactual explanations infer actionable changes, as Verma et al. (2024) said “[they] do not explicitly answer ‘why’ the model made a prediction; instead, they provide suggestions to achieve the desired outcome.” This suggests that the counterfactual effect of a comparative example is demonstrated by informing people what features have been changed, thereby altering model prediction. Only displaying a comparative example without highlighting the feature changes from the original example might not result in a perceivable and understandable counterfactual effect.
A.4. Rule-based Explanations
Rule-based explanations typically involve explaining model predictions by outlining one or several decision rules (Molnar, 2020). These rules might take the form of a simple “if…then…” statement or more complex logic using a combination of arithmetic operators, such as “AND…¿…=…”. These rules provide a clear and structured way for humans to understand how the model arrived at a specific decision.
Rule-based explanations can inform either local or global interpretations, depending on the specific techniques used to generate these rules (Aghaeipoor et al., 2023). For example, interpretable rule-based machine learning models, like decision rules, are trained on all instances to extract rule sets as model decision-making conditions (Molnar, 2020). In contrast, model-agnostic techniques like Anchors (Ribeiro et al., 2018) create rules that explain the relationship between inputs and outputs for instances close to a specific test instance. These nearby instances are created using a perturbation approach. By focusing on these neighboring instances, these rules only characterize how certain features of the test instance specifically affect the prediction, regardless of changes in the other feature values.
A.5. Concept-based Explanations
Concept-based explanations primarily present user-defined concepts embedded within the model’s latent space to explain its predictions in the form of natural language or a ranking of concept significance (Kim et al., 2018; Molnar, 2020). For example, a zebra (the image class predicted by AI) is generally identified by its black-and-white stripes (identified concepts).
Typically, concept-based explanations shed light on a model’s global behavior by recognizing and extracting shared semantic features for a particular class, such as colors or patterns, from the model embedding. More recently, local approaches are proposed, such as identifying the most influential regions of an image then associating these regions with broader, global concepts (Shukla et al., 2023) or using prompt-based image segmentation model to automatically detect concepts within an image (Sun et al., 2023b).
Appendix B Narrative Review Process and Corresponding Coding Scheme for Reviewed Articles
In this section, we detail the narrative review process, including search terms, databases, inclusion/exclusion criteria, paper grouping strategies, and coding schemes.
Our literature search was conducted in two main phases. The first phase, in Summer 2024, involved keyword searches across four digital scholarly databases: Web of Science, ACM Digital Library, ScienceDirect, and Taylor & Francis Online [^6] Web of Science: https://www.webofscience.com/, ACM Digital Library: https://dl.acm.org/, ScienceDirect: https://www.sciencedirect.com/, and Taylor & Francis Online: https://www.tandfonline.com/. . We searched the phrase “Explainable AI” AND “Design” combined with keywords “Framework” OR “Principle” in the title field. The second phase, in March–April 2025, focused primarily on the ACM Digital Library, targeting conferences closely related to XAI. Using the same search strategy, we targeted on major HCI conference proceedings from the past year that were not included in the first round, such as CSCW 2024, IUI 2025, and CHI 2025. Our initial screening yielded N = 273 relevant papers (Table 2).
Database Number of Articles
Web of Science Core Collection 92
ACM Digital Library 88
ScienceDirect 84
Taylor & Francis 9
Total 273
Publication date (1987 to 2025)
Table 2. Articles were identified through electronic database searches conducted in two phases. After scanning the abstracts, papers that did not mention “explainable AI” or “interpretable AI” were excluded. In total, 273 unique papers were included in the review corpus.
Category (N = 29) Representative Papers
User-centered (15) Understanding user needs:
Sanneman and Shah (2020), Liao et al. (2020), Xin He and Zhang (2023), Brennen (2020), Mucha et al. (2021), Hagras (2018), Vera Liao and Varshney (2021), Kaplan et al. (2024), Vera Liao et al. (2022)
Alignment between human cognition and model understanding:
Wang et al. (2019), Miller (2023b), Singh et al. (2021), Miller (2019), Bertrand et al. (2022), Dragoni et al. (2020)
Stakeholder-centered (5) Dhanorkar et al. (2021), Hoffman et al. (2021), Mohseni et al. (2021), Preece et al. (2018), Tomsett et al. (2018)
Socio-technical (5) Sokol and Flach (2020), Ehsan et al. (2021), Ehsan and Riedl (2020), Ehsan et al. (2023), Brandão et al. (2019)
Others (4) Forms of XAI interactions:
Bertrand et al. (2023), Chromik and Butz (2021), Xu et al. (2023)
Goal-oriented framework:
Suresh et al. (2021)
Table 3. Categorization of XAI Survey and Framework Papers
As described in Section 4, our review adopted a narrative approach rather than strictly adhering to the PRISMA protocol. Instead of reading all screened papers in detail, we selectively examined papers according to the strategies outlined below (Baumeister and Leary, 1997; Ferrari, 2015).
We organized relevant papers into these two categories: 1) survey and framework papers that examine the design process of building explainable AI systems, and 2) design-related or user evaluation papers that provide empirical evidence on the design effects of different XAI techniques or use cases. In this process, technical systematic papers that examine interpretable algorithms are excluded. Resulting survey papers are typically systematic review articles, either scoping XAI design practices (Bertrand et al., 2023; Kaplan et al., 2024; Mohseni et al., 2021) or proposing new framework for future XAI design (Wang et al., 2019; Miller, 2023a; Mohseni et al., 2021). Resulting framework papers include some that categorize high-level design guidance from interviews (Liao et al., 2020; Brennen, 2020) or proposed design methodology (Suresh et al., 2021; Dhanorkar et al., 2021). During this process, we also conducted a supplementary search to identify additional relevant articles using both backward and forward citation tracking.
After carefully reviewing the survey and framework paper corpus (N = 29), we categorized them into three focus areas: user-centered, stakeholder-centered, and socio-technical (see Table 3). This review helped us distill different emphases in XAI design and situate rhetorical XAI as a communicative act that complements and extends the goals emphasized in prior work (described in the Background section 2).
We then began annotating the collected design papers (N = 41) using our rhetorical framework (Tabel 1), which distinguishes between two goals, explaining how AI works and explaining why AI merits use, and three rhetorical appeals: logos, ethos, and pathos. As described in Section 4, a structured three-step was performed involving both open and axial coding. Table LABEL:tab:coding_scheme presents our coding scheme, which outlines the rhetorical elements and their illustrative mappings from the reviewed papers, along with data exemplars. Each piece of content was assigned a code indicating both the design intent it represents and the rhetorical appeal it conveys. The primary coding was conducted by the first author and subsequently validated by the second author.
Table 4. Reviewed XAI Design Papers: Mapping Design Strategies to Rhetorical Appeals
Paper (N=41) Mapping User Effect to Rhetorical Appeals
Alqaraawi et al. (2020) How AI Works: Isolated presentation
Logical Participant mentions of saliency maps features to understand the system:
p280: “…The rationale for this was that we aimed to compare how frequently participants mentioned concepts related features that saliency maps could potentially highlight…Our results show that when saliency maps were shown, participants predicted the outcome of the classifier significantly more accurately…However, even with the presence of saliency maps, success rates were still relatively low (60.7%)” p282-283: “…Furthermore, it is hard to get a quantifiable measure of the importance of individual features in an image. Again complexity increases if one attempts to quantify the importance of a feature on new images…Another reason why noticing Saliency-features does not necessarily facilitate a better understanding of a model is that general-attributes (e.g. colour, contrast) might influence the classification outcome…
Alvarado et al. (2022) How AI Works: Isolated presentation;
Why AI Merits Use: Expressive interaction
Credible Qualitative feedback about the increased ownership of using the system:
p20: “Center in User-Perceived Control and Awareness: …Moreover, participants described an enhanced sense of control and even ‘ownership’ over the recommendation system because TUIs gave them a ‘sense of touching, of power in their hands’.
Emotional Qualitative feedback about the enjoyability of using the system:
p17: “During the post-interview questionnaires, P15 supported these notions: ‘Playing around with them [RD and RP] will very likely improve the experience with the recommendations and, therefore, with the potential enjoyability of the movies/shows.’
Bansal et al. (2019a) How AI Works: Adaptive presentation
Logical Effects of explanations on human mental modal of the error boundary:
p4-5: “…in our controlled experiments we can vary the complexity of the error boundary and measure how it affects the accuracy of humans modeling the true error boundary….In the beginning, the worker makes more mistakes (more red crosses) because the mental model thus far is only partially correct. Eventually, the worker learns the correct model and successfully compensate for the AI (more red checks)…” p6: “…Since error boundary complexity increases with the number of conjunctions, we observe that a more parsimonious error boundary (i.e., a single conjunction) results in a higher team performance.
Bansal et al. (2019b) How AI Works: Adaptive presentation
Logical Effects of explanations on human mental modal of AI predictions:
p2431: “By analogy, we define that an update to an AI component is locally compatible with a user’s mental model if it does not introduce new errors and the user, even after the update, can safely trust the AI’s recommendations.” p2433: “…Figure 3a shows that, as the number of features increases, team performance decreases because it becomes harder to create a mental model…Figure 3b shows that as errors become more stochastic, it becomes harder to create a mental model, deteriorating team performance.
Bansal et al. (2021) How AI Works: Adaptive presentation; Why AI Merits Use: Technical Assurance
Logical Effects of explanations on human agreement on AI predictions:
p9: “Across the three datasets, Adaptive explanations successfully reduced the human’s tendency to blindly trust the AI (i.e., decreased agreement) when it was uncertain and more likely to be incorrect.
Qualitative feedback:
p11(table 3): “…Used AI as a starting references point…Double-checked after they made their own decisions…
Credible Qualitative feedback:
p11(table 3): “…Explicitly acknowledged they used the confidence…Was more likely to accept AI above the threshold
Ben David et al. (2021) How AI Works: Isolated presentation, Protective framing;
Why AI Merits Use: Explanation as recommenders
Logical Effects of explanations on self-reported readiness to adopt and willingness to pay:
p394-5: “On the very first day when the model was introduced (day 3) we see an average of 60% adoption when averaging over all participants…Following the model failure on day 7, adoption plummeted, returning approximately to the initial level with 61% on day 8, this was followed by a slow recovery up until day 10 (66%)…Finally, when participants were told advice will no longer be free, and were asked about willingness to pay, we see an overall decrease in adoption on the final day (77%).
Emotional Algorithmic aversion:
p397: “…the two more elaborate explanations, the Performance-Based (P-value = 0.0445, t = -2.0225) and the Feature-Based (P-value =0.0498, t = -1.9741) prove to be more resistant to the effect of algorithm error…When we explore the adoption recovery…we find that the Performance-Based explanation had a marginally significant (P-value = 0.0633, t = 1.8691) tendency to recover better than the global explanation (12% compared to only 2% for the Feature-Based and the Global Explanation, respectively)…
Bernardo and Seva (2023) How AI Works: Isolated presentation, Explanation affect, Emotional cues
Logical Effects of explanations on self-reported usefulness:
p15: “Example-based explanation increases interestingly surprised, trusting emotions, and perceived usefulness, while feature-and rule-based explanation decreases them. Logic robotic increases fearfully dismayed plus anxiously suspicious emotions and decreases perceived usefulness, while humanized communication functions the opposite. The presence of supplementary information decreases fearfully dismayed emotions, while the absence of it increases the effect.
Credible Effects of explanations on self-reported trust and reliance (see above summary findings)
Emotional Participant AI anxiety and incidental emotion influence user perceived trust of explanations:
p16: “First, when fearfully dismayed emotions were felt, users with high AI anxiety experienced lower perceived trust than those who have low AI anxiety.”
Bove et al. (2022) How AI Works: Mechanistic presentation
Logical Effects of explanations on self-reported system understanding:
p814: “We propose a questionnaire approach…Each item in the questionnaire is a statement, for which users can either answer ”true”, ”false” or ”I don’t know”. We design three types of questions to capture different components of user understanding: (i) explanations’ scope…(ii) explanations’ effects…(iii) explanations’ locality” p816: “Contextualization improves objective understanding…Exploration does not have a significant impact…The interaction of contextualization and exploration does not have a significant impact neither.
Credible Effects of explanations on self-reported satisfaction:
p816: “Contextualization significantly improves users satisfaction…contextualization parameter is greater than the claimed value and conclude that contextualization significantly improves non-expert users’ satisfaction…the exploration parameter is greater than the claimed value and conclude that exploration significantly improves non-expert users’ satisfaction
Cai et al. (2019a) How AI Works: Isolated presentation
Logical Effects of explanations on self-reported system understanding, capability, and benevolence:
p3-4“When the algorithm failed to recognize the drawing, users who received normative explanations felt they had a better understanding of the system…Likewise, when the algorithm did not recognize the drawing, users also rated system capability higher if they had seen a normative explanation…In addition, we found a small but significant main effect of comparative explanations on system benevolence. Users who saw comparative explanations perceived the system to be more benevolent.
Cai et al. (2019b) How AI Works: Mechanistic presentation;
Why AI Merits Use: Explanation as retriever
Logical Effects of explainable interactive tool on self-reported diagnostic utility, future use, and overall preference:
P8(figure 5): “Compared to a conventional interface, SMILY had greater diagnostic utility and required lower effort. Users also indicated having greater trust in SMILY, felt SMILY offered better mental support, and felt they were more likely to use it in clinical practice.
Qualitative feedback:
p10-11: “Tracking the Likelihood of a Decision Hypothesis. During the study, users described ways in which iteratively refining search results helped them track the likelihood of a hypothesis…Generating New Ideas. The refinement tools also provided decision-making support by helping users reflect on their thought process and generate new ideas…Reducing Problem Complexity. Reference images containing a mixture of features were often cognitively difficult to analyze…
Credible Effects of explanations on self-reported trust in the system:
p9: “In sum, users were able to retrieve more diagnostically relevant information with less effort using SMILY, had greater trust in the system, and felt they were more likely to use it in clinical practice…
Cau and Spano (2025) How AI Works: Mechanistic presentation; Why AI Merits Use: Technical assurance
Logical Effects of explanations on accuracy and overreliance:
p1449: “…Consequently, we fail to reject the null hypotheses for H1a and H1b, concluding that the explanation type does not affect users’ accuracy and overreliance on AI…
Self-reported understanding and usefulness of explanations:
p1450-1451: “…15 participants perceived Local Model Explanations (LME) as fast and time-saving support for accepting/rejecting an application…12 participants highlighted the difficulty in understanding why the AI assigned specific weights to attributes and the heavy emphasis on the computer skills attribute…10 participants were overwhelmed by excessive information and suggested improving the visualization of the data distribution module.
Qualitative feedback:
p1451: “…10 participants were overwhelmed by excessive information and suggested improving the visualization of the data distribution module…
Credible Self-reported user confidence and trustworthiness of explanations to use AI:
p1453: “…our exploratory analysis revealed an association between requesting on-demand assistance and lower user confidence. This suggests that accessing data or AI support can undermine confidence, regardless of the type of assistance provided. However, it remains unclear whether explanations reduced confidence or if participants requested assistance due to feeling less confident.
Qualitative feedback:
p1451: “…The Data-Centric Explanations (DCE) were perceived by 18 participants as trustworthy, giving clear guidelines for the job application requirements…
Chen et al. (2023b) How AI Works: Isolated presentation
Logical Effect of explanations on decision-making accuracy and reliance:
p15: “…Across tasks, example-based explanations achieved complementary decision performance whereas feature-based explanations did not…” Qualitative feedback:
P13: “With feature-based explanations, participants applied intuition about features to determine if they agreed or disagreed with the model’s reasoning…example-based explanations can provide additional context and support inductive reasoning
Combined summary:
Example-based explanations led to less disruption of people’s natural intuition about the outcome…Example-based explanations better supported recognizing or even learning new features because the similar examples and their ground truth labels brought in additional context that could promote inductive reasoning…Example-based explanations provided strong signals of unreliability, particularly when the AI system predicts both examples incorrectly.
Díaz-Rodríguez and Pisoni (2020) Why AI Merits Use:
Authority framing, Cultural relatedness
Credible Qualitative feedback:
p319: “…Does embedding of expert/domain knowledge into DL models help explain such models? Can XAI help encode such prior knowledge? Juan Jesus Pleguezuelos, History teacher and podcast author of Art History for entrance exams to University: The challenge I pose is to make others see an historical image only through words. It is clear that this requires an exhaustive description of the masterpiece, but you should also try to make others feel the latent soul in it, and decipher the intention of the author…
Emotional Qualitative feedback:
p319: “…And if you could also convey the emotion that this work is able to cause, it can be that words may be more than enough to make a listener understand an artistic work that he is not seeing in that moment…
Ebermann et al. (2023) How AI Works: Isolated presentation;
Why AI Merits Use: Protective framing
Logical Self-reported evaluation of AI support:
p1815: “This suggests that cognitive misfit reduces the subjective support. However, in contrast to our expectations, the evaluation improved from T2 to T3 in the case of decision situations with cognitive misfit…value on the variable ‘evaluation of AI’s support’ was higher at the third time point of measure…The ‘evaluation of AI’s support’ is indeed lower when cognitive misfit occurs.
Credible Self-reported willingness to use AI:
p1816: “In the case of cognitive misfit resulting from different decisions and corresponding explanations of the AI and the user, users do not have significantly less willingness to use the AI again in comparison to cases where cognitive fit occurs.
Emotional Self-reported mood score:
p1815: “In line with our expectation…where the decision and explanation of the AI is known and cognitive dissonance might occur, that the values of the variable ‘mood’ are significantly lower in the case of participants in decision situations with cognitive misfit…
Ehsan et al. (2021) How AI Works: Isolated presentation; Why AI Merits Use: Social endorsement
Credible Qualitative themes:
Supported Effects…Understanding organizational norms and values, social validation, transitive trust, and temporal relevance…
Ehsan et al. (2024a) How AI Works: Isolated presentation, Numerical association
Logical Self-reported intelligence and understandability:
Participants with AI background also viewed numbers as potentially actionable even when their meaning was unclear…Unable to access their meaning, the non-AI group associated numbers with the presence of a higher, more intelligent expression…both AI and non-AI groups found unanticipated explanatory value in AD’s declarative statements and NR’s numerical representations…
Credible Self-reported confidence, friendliness, and second chance use:
Participants in both groups had unwarranted faith in numbers. However, their extent and reasons for doing so were different…The mere presence of numbers was associated with an algorithmic thinking process in the robot even when the meaning of numbers was unclear…The AI group seems to have followed heuristic reasoning that associates mathematical representations with logic and intelligence…
Evans et al. (2022) How AI Works: Isolated presentation, Graphical synergy;
Why AI Merits: Explanation as recommenders
Logical Self-report on “explanations are understandable” and “explanations are useful for decision-making”:
p286: “Explanation examples are presented in descending order of total positive responses (slightly agree to strongly agree)to the statement: ‘The explanation provides me with valuable information for my work.’ Trust scores were the most widely accepted explanation by this metric, while the explanation receiving the most positive median response was the Counterfactuals (One-axis) example.
Qualitative feedback:
P288-289: “…Two saliency map examples difficult to understand, distracting and/or confusing (p1, 2, 5), to trivially understandable but not at all useful (p6)…The risk for positive confirmation bias with such [saliency] explanations was pointed out: “[It] is hiding a lot of information. But it’s very tempting, because it’s very simple” (p1)
P289: “These [concept] examples were widely considered intuitively understandable and informative about the factors important for the AI output, albeit with limited value ascribed to this information (p1, 2, 4, 6)…Criticisms of the explanation were the lack of precision and/or descriptiveness (p1–3), that having to read text is generally undesirable in the pathologist workflow (p3, 4)
P290: “Pathologists were almost unanimous in expressing that these [counterfactuals] explanations ‘[helped them] understand what the algorithm is looking for’ (p1) All participants drew some concrete conclusions about the factors important to the AI solution, finding it self-explanatory that staining was the most important factor distinguishing positive nuclei from those marked negative, with some identifying other important factors such as shape, size of nuclei, particularly in separating the positively annotated from the unclassified nuclei (in the two-axis example) (p1–3).
Credible Qualitative feedback:
p290: “The labeling of annotations as high- or low-confidence was generally regarded as being intuitively understandable. Aside from identifying the AI solution’s confidence in its annotations, many participants also inferred from this the factors that may have been important to the solution.
Freire et al. (2023) How AI Works: Isolated presentation, Emotional cues
Credible Effects of emoji and explanations on human-AI agreement:
The explainable AI (XAI)-emoji and XAI conditions have higher agreement scores than the emoji and control (no intervention) conditions.
Gaole et al. (2025) How AI Works: Adaptive and Mechanistic presentation
Logical Effects of explanations on user objective feature understanding and appropriate reliance based on human-AI agreement:
p914-915: “We found thatbLLM Agent condition achieved significantly worse objective feature understanding than the Dashboard, CXAI, and ECXAI conditions…In our study, we found that interactive XAI interfaces can have a negative impact of increasing over-reliance on the AI system.
Credible Effects of explanations on self-reported trust of the AI system and user confidence:
p915: “With post-hoc Tukey’s HSD test, we found that participants who received XAI showed significantly higher trust in Understandability/Predictability (i.e., Control ¡ Dashboard, CXAI, ECXAI).”
Emotional Effects of explanations on self-reported engagement of using the system (No direct report in the result)
Kenny et al. (2021) How AI Works: Isolated presentation;
Why AI Merits Use: Explanation as retriever
Logical Effects of explanation on correctness:
p17: “…explanations seem to mainly have an effect when things go wrong, when errors arise and outputs diverge from what a user expects/desires…
Self-reported reasonableness of explanations:
p21: “…this system-level measure of correctness (like the trust measure) shows that the item-level effects for Explanation do not persist into a system-level evaluation of correctness. The Error-Rate’s variable was also found to impact overall correctness and overall reasonableness; see Table 2), showing people are sensitive to changes in error-rates from 4% to 12% and 28%.
Credible Self-reported trust and satisfaction:
p21: “…Specifically, the Explanation-Present group rated Question 1 (“From the explanation, I understand how the program works”), Question 3 (“This explanation of how the program works has sufficient detail”), and Question 5 (“This explanation of how the program works tells me how to use it”) significantly higher than the Explanation-Absent group, all p ¡ .05…
Kim et al. (2024) How AI Works: Mechanistic presentation;
Why AI Merits Use: Explanation as recommenders
Logical Qualitative feedback:
p7-8: “…clinicians found the interfaces helpful in making informed decisions and determining the order of importance of different aspects…To enhance their experience, patients suggested some improvements, including the provision of visual cues to facilitate a more intuitive understanding of their current status. One patient mentioned the following regarding the color-coding scheme…The clinicians expressed a strong desire to identify the types and causes of errors made by their patients…
Kocielnik et al. (2019) How AI Works: Isolated presentation;
Why AI Merits Use: Expectation management
Logical Self-reported understanding of AI model, perceived control over system behaviors:
p8: “Comparing conditions with and without Explanation revealed a significant positive impact…with higher average level of perceived understanding (mean of both understanding questions) for participants who saw the Explanation than for those who did not…We therefore consider H2.3 supported as the Control Slider intervention significantly increased participants’ perceived level of control over the Scheduling Assistant’s behavior.
Emotional Expected satisfaction of system performance:
p10: Participants in the High Recall version of the system reported no significant difference in satisfaction between the Baseline and Pure-techniques. At the same time, in the High Precision version, participants exposed to Pure-techniques reported significantly higher satisfaction than those in the Baseline…
Liu et al. (2021) How AI Works: Isolated presentation with interactivity
Logical Effects of explanations on appropriate trust (reliance):
p22: “In general, we do not find significant impact from interactive explanations with respect to the performance of human-AI team or human agreement with wrong AI predictions, compared to static explanations. However, humans are more likely to find AI assistance useful with interactive explanations than static explanations in ICPSR and COMPAS, but not in BIOS…
Self-reported real-time assistance useful:
p25: In COMPAS, Interactive/Interactive achieves a significantly higher human perception of real-time assistance usefulness than both Static/Static and Interactive/Static…Perception of Interactive/Static is also significantly higher than that of Static/Static. We find similar results in ICPSR except that the difference between Static/Static and Interactive/Static is not significant…
Self-reported identification of important features:
p25: “These results suggest that participants with interaction are more likely to fixate on demographic features and potentially reinforce human biases,11 but are less likely to identify computationally important features in ICPSR and COMPAS.
Ma et al. (2023) How AI Works: Adaptive presentation; Why AI Merits Use: Technical assurance
Logical Effects of explanation on human-AI agreement (appropriate trust):
Figure 6 shows the human-AI agreement in different human-AI CL situations under three conditions. Results showed that all three conditions made people’s agreement with AI significantly higher when AI’s CL was higher than that when the human’s CL was higher…the human-AI agreement in AI Confidence was significantly higher than Adaptive Workflow , and Adaptive Recommendation, and it was marginally higher than Direct Display. Note that when AI is wrong, a lower agreement with AI is better. These results suggest humans’ less over-trust in AI in our proposed CL exploitation conditions…However, when the AI’s recommendation was correct, we did not observe significant differences in human-AI agreement between the three proposed conditions and AI Confidence baseline.
Qualitative feedback:
Participants in Adaptive Workflow and Adaptive Recommendation were forced to think independently.
Credible Qualitative feedback of correctness likelihood:
Some participants doubted the displayed CL information and ignored it in the decision-making process…Most participants referred to or were influenced by the dis- played CL or CL-based adaptation
Nguyen et al. (2018) How AI Works: Isolated presentation with interactivity
Logical Effect of presenting AI predictions on fact-checking accuracy:
p194: “We observe that there are larger changes for errors made by participants in group Control: those who have not seen our stance predictions change their answers more than those who have…This roughly corresponds to the errors by the veracity classifier, showing again that system predictions can both help users (when correct) or lead users to errors that reflect model fallibility or biases implicit in training data.
Qualitative feedback:
p196: “It was very hard to understand. It seems on one task, I was 100% sure it was true and I was told it was false, I even followed links to verify the sources
Emotional Qualitative feedback:
p196: “…thought it was really cool! I’d enjoy playing with this more if it wasn’t during my work time…
Okolo et al. (2024) How AI Works: Isolated presentation, Cultural relateness, Emotional cues
Logical Qualitative feedback:
p11: “While these interpretations made sense to the CHWs, they were not what the XAI methods intended to convey (they intended to show the original image for reference and the annotated image for explainability, with no storyline among the images). Only a few CHWs could partially understand the explanations with the help of higher-level features such as colors, images, or shapes” p17: “This interpretation was incorrect because the underlying model relied on specific bodily regions (in alignment with the Kramer’s Rule chart in Table 1) to provide a severity prediction, not on the number of yellow boxes.
Credible Qualitative feedback:
p13: “P04 and P05 emphasized that their patients trust them more when they use visual aids. They gave the example of paper-based visual aids they carry to educate pregnant women about when to seek urgent care. When we asked if CHWs would prefer similar paper-based visual aids for jaundice, P04 and P05 preferred the AI-driven app over paper-based visual aids, stating that the app could diagnose the severity of diseases, which the paper-based aids could not.
Emotional Qualitative feedback:
p8: “They were worried that the red color might confuse the participants since it is often used in public health messaging to imply ‘danger’ in the medical sense.” p14: “CHWs often relied on their prior experience with colors in visual aids in the domain of Public Health to interpret what the colors in the XAI visualizations might mean. They had strong color associations, believing that ‘green means safe, yellow means warning, red means danger and gray/blue means lack of blood.’ These color associations largely influenced how they interpreted the visualizations depicting a jaundiced baby.
Okoso et al. (2025) How AI Works: Authority framing, Emotional cues
Credible Self-reported acceptance of AI advice:
In the Formal, Authoritative, and Humorous groups, an upward trend was observed in d tone u, indicating that participants tended to increase their scores. Although no significant overall effect of tone was observed, the tone could possibly affect participants differently depending on specific user attributes.
Qualitative feedback:
Among the 135 participants who noticed a change in tone and provided relevant responses, 59 participants said it did not affect their decisions, whereas 55 participants said it did. Some participants felt that an informal tone was inappropriate for a serious topic, such as recidivism risk, whereas others found the tone easier to understand.
Emotional Qualitative feedback:
A total of 110 participants (37%) expressed either positive or negative opinions about the explanations but ultimately stated that it did not influence their decision. Most felt that the explanations did not provide sufficient information relevant to their needs…those movies based solely on genre and did not consider the explanation, whereas others remarked that it felt unnatural for an AI, which supposedly lacks emotions, to promote a movie
Schemmer et al. (2023) How AI Works: Isolated presentation
Logical Effects of explanation on appropriate reliance:
p418: “In the XAI condition, we can observe a significant increase in relative AI reliance (PAIR) from 29.59% to 38.87% while the (relative self-reliance) RSR does not change significantly (71.87%) for the control condition and 69.45% for the feature importance condition. This means explanations of AI decisions can reduce the share of under-reliance. It is important to highlight that PAIR is not increased simply by relying more often on AI advice, as this would have also reduced the RSR significantly.
Schuff et al. (2022) How AI Works: Isolated presentation
Logical Self-reported importance of features reflect bias of understanding explanations:
p619: “Overall, three studies confirmed the presumably biasing influence of word length, (pairs of) two studies respectively confirmed the effect of sentence length, display index and saliency rank, and one study (each) found significant effects of word position, sentiment polarity, word frequency, capitalization and dependency relation
Sun et al. (2023a) How AI Works: Isolated presentation; Why AI Merits: Explanation as retriever
Logical Improve model performance by using local explanation-based attention steering and evaluate the model vulnerability:
p19-20: “S2 results suggest that (1) the workflow of the local explanation-based attention steering provided a diverse perspective in diagnosing model vulnerability, (2) the direct steering design helped the process of model revision straightforward, and (3) every participant enjoyed improved key model performance measures…The behavioral data we collected shows that all participants generated the model that outperforms (1) its model accuracy, (2) the overlap between the model’s focus and the relevant object types (IoU), and (3) the proportion of reasonable attention out of all images in a test set…After completing the user studies, the majority of users strongly agree that adjusting local explanations can effectively improve model performance.
Qualitative feedback:
p20: “During interviews, all participants shared their positive impressions about the effectiveness of attention adjustment in improving model accuracy, which is the primary objective of conducting model fine-tuning. They also confirmed that the impact of contextual bias was reduced as attention quality increased by attention steering” p22: “Existing issues: insufficient description about reasonability matrix, Ill-formed local explanation visualization…
Suresh et al. (2022) How AI Works: Isolated presentation with interactivity;
Why AI Merits Use: Explanation as retriever
Logical Effects of explanation on human-AI agreement:
p776: We recorded the percent of cases in which participants agreed with the model (versus when they disagreed or were not sure). For cases in which the prediction was correct, the agreement rate was similar across conditions; however, when the prediction was incorrect, we found that participants were less likely to accept the model’s prediction when they were using the NN interface, with or without the input editor. Often in these cases, they did not explicitly “disagree” with the model, but wanted additional information about the signal and/or patient before committing to an answer.
Qualitative feedback:
p776: “Nearest neighbors enable reasoning with clinically-relevant concepts…In contrast, with the baseline condition, participants often had difficulty extracting higher-level, clinically-relevant concepts from the feature importance visualization…Nearest neighbors help characterize uncertainty and incorporate it into decision-making
van der Waa et al. (2021) How AI Works: Isolated presentation; Why AI Merits Use: Explanation as recommender
Logical Identification of model decisive factor and advice:
p15: “First, rule-based explanations indeed seem to allow par- ticipants to more accurately identify the factor from a situation that was decisive in the system’s advice. However, rule-based nor example-based explanation allowed participants to learn to predict system behavior.
Effects of explanations on perceived system understanding and prediction of correctness:
p15: “The rule-based explanations however, did cause to participants to think that they better understood the system compared to example-based and no explanations. The example-based explanations only showed a small and insignificant increase in perceived system understanding.
Vasconcelos et al. (2023) How AI Works: Isolated presentation
Logical Effects of explanations on overreliance and the influence of task difficulty:
p14: “an increase in task difficulty increases the cost of following the strategy to engage with the task, as it generally takes more cognitive effort for people to either (a) ensure that the AI has gotten the answer correct when provided with only predictions or (b) complete the task alone. However, the strategy to rely on the AI does not incur the same costs, as people are not engaging with how hard the task is.” p21: “the salient explanation condition has less overreliance than all other conditions (prediction, highlight explanation, written explanation, and incomplete explanation…We also find that the incomplete explanation condition has less overreliance than the prediction only and written explanation conditions, more overreliance than the salient explanation condition and finds no difference in overreliance in the highlight explanation condition.).”
Wambsganss et al. (2020) How AI Works: Adaptive and Mechanistic presentation;
Why AI Merits Use: Explanation as recommender
Logical Self-reported intention to use, perceived usefulness and ease of use:
p9: “the mean value of intention to use of participants using AL as a writing tool was 2.33 (SD= 0.58). These values are significantly better than the results of the alternative scripting approach. For perceived usefulness we observed a mean value of 3.2 (SD= 1.12) and for perceived ease of use the value was 2.83 (SD= 1.08) for participants from the control group. The mean value for the intention to use was 3.5 (SD= 1.13). The results clearly show that the participants of our experiment rated the acceptance of AL as an adaptive feedback tool positively compared to the usage of the alternative application. The statistical significance was also proven in a double-sided t-test for all three constructs (see table 5).
Qualitative feedback:
p9: “The general attitude for AL was very positive. Especially the fast and direct feedback, the graph-like visualization of the argumentation structure and the summarizing scores were mentioned quite often. However, sometimes AL was not correctly classifying claims and premises, which users suggest to improve.
Wang et al. (2019) How AI Works: Mechanistic presentation;
Why AI Merits Use: Explanation as recommender
Logical Qualitative feedback:
p10-11: “Finding Alternative Hypotheses. We found many instances of users seeking alternative hypotheses with or without various types of XAI…P4 found counterfactual reasoning useful even though one cannot change a patient’s history, ’because it’s a matter of possibility. It’s how you rank your differentials…
Credible Qualitative feedback:
p11: “Lack of Trust: Verify with Supplementary, Situational, Source Data. A lack of trust drove users to want to verify the AI’s decision by looking at raw data, in this case, patient case and physical exam notes, or recorded data…without the context of the patient, [she] can’t diagnose based on the vitals itself. So [she] wouldn’t really trust
Wang and Yin (2021) How AI Works: Isolated presentation
Logical Effects of explanations on human understanding of model behaviors and calibrated trust:
p325-326: “The only positive effect of model explanation that we have consistently observed across different decision making tasks is that feature importance explanations increase people’s objective understanding of an AI model, while feature contribution explanations increase people’s subjective understanding of an AI model…feature contribution explanation seems to be able to satisfy more desiderata of AI explanations when people have some domain expertise in the decision making task…For the two example-based explanations in our study, we found minimal evidence on their capability to support trust calibration.
Wang and Yin (2023) How AI Works: Adaptive presentation
Logical Effects of explanations and model updates on perceived system understanding:
p8: “We find that participants’ perceived change of explanations increased as the the explanations of the updated model in Phase 2 became more dissimilar from those of the old model used in Phase 1. In other words, participants in our experiment could perceive the change in model explanations brought up by a model update.
Effects of explanations and model updates on objective just (frequency of accepting model decisions):
p8: “participants’ objective trust gain and subjective trust gain in the AI model from Phase 1 to Phase 2, both across all participants and within subgroups of participants with different levels of understanding of the AI model. However, we find that neither participants’ objective trust nor their subjective trust seems to be affected by the similarity level of model explanations between Phase 1 and Phase 2.
Credible Effects of explanations and model updates on subjective trust gain and satisfaction gain:
See above excerpt about objective trust; p8: “Figure 3(d) further shows participants’ subjective satisfaction gain from Phase 1 to Phase 2, conditioned on their understanding score. Still, participants did not seem to significantly change their satisfaction with the AI model as the similarity of model explanations before and after the update varied.
Wang et al. (2023b) How AI Works: Mechanistic presentation, Emotional cues; Why AI Merits Use: Explanation as recommender
Logical Effects of counterfactual explanations on several tool useful measures: ‘easy to use,’ ‘easy to understand,’ ‘help understand model,’ ‘help find ways to improve plans,’ and ‘will apply loan again’:
p11: “(A) Participants thought GAM Coach was relatively easy and enjoyable to use, and the tool helped them identify actions to obtain a preferred ML decision. (B) All interaction techniques, especially experimenting with hypothetical values, were rated favorably.
Qualitative feedback:
p9: Three main reasons that participants reported choosing plans were that the plans were (1) controllable, (2) requiring small changes or less compromise, or (3) beneficial for life in general…Participants’ explanations for not choosing a plan mostly complemented the reasons for choosing a plan…Preference configuration is helpful…Participants’ explanations for not choosing a plan mostly complemented the reasons for choosing a plan. Some participants also skipped plans because they were puzzled by counterintuitive suggestions, did not understand the suggestions, or just wanted to see more alternatives.
Emotional Effects of explanations on enjoyable to use (See previous excerpt on tool usefulness in logical)
Yang et al. (2020) How AI Works: Isolated presentation
Logical Effects of explanation and spatial layout on user appropriate trust of AI predictions:
p195: “The results support that all the visual explanations largely increase appropriate trust, reduce overtrust, and help users gain more confidence in their decisions…image-based explanations help correct undertrust, but the effect of rose-based explanations on undertrust is inconclusive…For the image-based explanations, all the three spatial layouts lead to a similar level of users’ trust; appropriate trust, overtrust, and undertrust are very similar and display few differences among the three spatial layouts
Effects of explanation on perceived helpfulness:
Helpfulness was measured but not directly reported in results
Credible Effects of explanation on self-confidence, trust meter before and after AI feedback:
p196-197: “Before seeing the feed- back, participants increase the trust meter for a correct recommendation and decrease the trust meter for an incorrect one; having an explanation shows small positive effects on trust direction…Image-based explanations outperform rose-based explanations because they increase appropriate trust, decrease overtrust and undertrust, improve self-confidence, and show more usability
Zhang et al. (2020) How AI Works: Isolated presentation
Logical Effects of explanation on human-AI agreement:
The baseline condition (green) and the explanation condition (blue) had similar agreement percentages, while the with-confidence condition (orange) had higher percentage when the confidence level was above 70%.